### v2

# Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning

Abstract
Knowledge Distillation (KD) focuses on using a teacher model to improve a student model. Traditionally, KD is studied in an offline fashion, where a training dataset is available before learning. In this work, we introduce the problem of Continual Distillation Learning (CDL) that considers KD in the Continual Learning (CL) setup. A teacher model and a student model need to learn a sequence of tasks, and the knowledge of the teacher model will be distilled to the student to improve the student model in an online fashion. The CDL problem is valuable to study since for prompt-based continual learning methods, using a larger vision transformer (ViT) leads to better performance in continual learning. Distilling the knowledge from a large ViT to a small ViT can improve inference efficiency for prompt-based CL models. To this end, we conducted experiments to study the CDL problem with three prompt-based CL models, i.e., L2P, DualPrompt and CODA-Prompt, where we utilized logit distillation, feature distillation and prompt distillation for knowledge distillation from a teacher model to a student model. Our findings of this study can serve as baselines for future CDL work. [^1] Project website https://irvlutd.github.io/CDL/
1 Introduction
Continual Learning (CL) [30] designs models that can continuously learn new tasks without forgetting previously learned tasks. In class-incremental continual learning, data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training. In testing, the model will be tested on all the seen classes. Therefore, a good CL model should learn new classes without forgetting.
Traditionally, the class-incremental setting is tackled using a rehearsal buffer [24, 17, 21, 3], which stores training examples from previous classes. During training, data from new tasks and data from the rehearsal buffer are used to update the model weights. With the help of the rehearsal buffer, the model can continually refresh its memory of previous classes. However, due to the size limit of the rehearsal buffer, these models still suffer from the catastrophic forgetting problem [22].
Recently, a new paradigm for CL based on learning for prompt receives more attentions [34, 33, 28, 31]. Instead of using rehearsal buffers, these methods focus on learning prompts for pre-trained large models such as vision transformers for continual learning. For example, L2P [34] learns a prompt pool for a set of tasks using pre-trained ViTs [6], where different tasks learn to use different prompts in the prompt pool. DualPrompt [33] learns two types of prompts, i.e., a global prompt for all tasks and task-specific prompts. By leveraging the pre-trained ViT models, these models achieve better performance compared to traditional rehearsal buffer-based methods.
We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance. For example, ViT large is better than ViT base, and ViT base is better than ViT small, etc. Therefore, it is appealing to use large models in prompt-based CL. However, larger models introduce more computation during inference. To address this problem, we study the problem of Knowledge Distillation (KD) from large to small models in the continual learning setting. We name this problem Continual Distillation Learning (CDL).
Figure 1: Comparison between three different knowledge distillation scenarios. (a) Traditional KD with an offline dataset. (b) KD in rehearsal-based CL models. (c) The CDL problem introduced in this work.
The CDL problem is different from the previous knowledge distillation setup. We illustrate these differences in Fig. 1. First, in the traditional KD setup, an offline dataset is used for knowledge distillation (Fig. 1(a)). A teacher model is first trained using the dataset, and then its knowledge is distilled to a student model using the same dataset. The setup is not the continual learning setup. Second, KD is also used in the rehearsal-based CL methods such as iCaRL [24] (Fig. 1(b)). These methods maintain a rehearsal buffer of previous data. When updating the model parameters for the current task, KD ensures that the prediction on the previous data should be similar to the prediction of the old model. In this case, the old model before updating is the teacher model. Lastly, the CDL problem we study in this work is illustrated in Fig. 1(c). Given a new task in continual learning, a teacher model is first updated. Then a student model is updated based on the data of the new task and the teacher model. Consequently, a better student model can be trained by knowledge distillation from the teacher model.
To study the CDL problem, we have designed a prompt-based continual learning method that consists of a teacher model and a student model. In our method, we jointly learn the teacher model and the student model. Both models are prompt-based CL models, where the backbones are frozen and the prompts are learnable. When training data for new tasks come, the teacher model is first updated as other prompt-based CL models in the literature, and then the student model is updated.
Our method is flexible to be used with different prompt-based CL models and different mechanisms for knowledge distillation. Therefore, we conducted an empirical study of the CDL problem using three prompt-based CL models, i.e., L2P [34], DualPrompt [33] and CODA-Prompt [28]. For each model, we considered different teacher-student configurations. For example, using ViT large as the teacher model and ViT base as the student model. In addition, we studied the effect of three types of knowledge distillation mechanism in the CDL setup, i.e. logit distillation, feature distillation, and prompt distillation. We found that prompt distillation works the best, due to the prompt learning nature of these models.
In summary, our contributions are as follows.
  • We introduce the problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
  • We designed a method for CDL that can be used with different prompt-based CL models and different knowledge distillation mechanism.
  • We conducted extensive experiments and analyses of three prompt-based CL models with three types of KD mechanism on two commonly-used benchmarks for continual learning.
2 Related Work
Continual Learning. The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting  [22] in the case of incremental tasks. Different types of continual learning methods have been proposed in the literature [32]. For example, [24, 17] utilize function regularization to help the loss function. There are also weight regularization methods [15, 38, 20, 25, 27] that selectively constrain changes in network parameters and impose penalties on changes in each parameter based on its contribution. Architecture-based approaches [37, 13, 35, 7, 23] mainly focus on constructing a special model for continual learning. Replay-based approaches [24, 17, 12, 21, 1, 3, 2] use the memory buffer to store and replay old task data in learning the current task. Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [9, 8, 5]. The aforementioned methods do not leverage pre-trained networks. They usually learn the tasks from scratch.
Prompt-based Continual Learning. Inspired by the use of prompts in natural language processing [19], prompt-based continual learning methods employ large Vision Transformer (ViT) models [6]. These methods freeze the pre-trained backbone and shift learning to the prompts of the ViTs. For example, L2P [34] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool to select the corresponding prompts. Based on this idea, DualPrompt [33] introduces the concept of using a general prompt and an expert prompt. CODA-Prompt [28] abandons the key-prompt pair selection idea. It uses a weighted sum of the prompt components to obtain the final prompt. In prompt-based methods, the size of the backbone matters. Larger backbone models achieve better performance. This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to speed up inference time.
Knowledge Distillation (KD). The purpose of knowledge distillation is to leverage large models to assist smaller models in fully exploiting their potential. The larger model, termed the teacher, aids in training the smaller model, referred to as the student. According to the distillation location, traditional methods can be divided into logit distillation [11, 39] and feature distillation [18, 26, 4]. With the recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [29, 36]. However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
3 Prerequisites
3.1 L2P
L2P [34] and DualPrompt [33] use a pre-trained ViT model as the backbone and keep the backbone frozen. Both methods use a prompt pool to encode information about tasks: $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}$, where $P_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompts. In order to select prompts for different tasks, each prompt is associated with a learnable key $\{(\mathbf{k}_{1},P_{1}),(\mathbf{k}_{2},P_{2}),\ldots,(\mathbf{k}_{M},P_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ and $\mathbf{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys.
Given an input image $\mathbf{x}$, a query function $q(\cdot)$ is used to encode the input image. The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(x),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function. Then the top $k$ prompts from the prompt pool $\mathbf{P}$ based on the establishment of key-value pairs are selected. These prompts are used in the ViT model to classify the input. In this way, given an input-label pair $(\mathbf{x},y)$, L2P optimizes the trainable prompt keys and values, and the final classification layer. The training loss function is defined as
$\min_{\mathbf{P},\mathbf{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (1)
where $g_{\phi}$ is the classifier, $f_{b}$ is the pre-trained ViT backbone, $\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification. The final class token of the ViT is used for the classifier. Therefore, the first part in Eq. (1) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathbf{P}$, key set $\mathbf{K}$ and the classifier parameter $\phi$. The second term minimizes the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
3.2 DualPrompt
The DualPrompt method [33] supplements insertable prompts based on the L2P method. The prompts are divided into G-Prompt (General) and E-Prompt (Expert). The shared G-Prompt among all tasks and the corresponding E-Prompt are attached to multiple multi-head self-attention (MSA) layers of the pre-trained transformer. The prompts in both the E-Prompt and L2P are identical, primarily serving to distinguish between different tasks. In contrast, the G-Prompt serves as a shared prompt across different tasks, mainly representing the shared information between tasks. In training, DualPrompt optimizes the G-prompt and E-prompt jointly. Please see [33] for details.
3.3 CODA-Prompt
Based on DualPrompt, CODA-Prompt [28] desires an approach that is end-to-end differentiable. Without selecting the top-K key-value pairs in the prompt pool, CODA-Prompt builds the prompt components: $\mathbf{p}=\sum_{m}\alpha_{m}P_{m}$, where $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}\subseteq\mathbb{R}^{L_{p}\times D\times M}$ is the set of prompt components, $M$ is the length of the set. $\alpha_{m}$ is the weight for the component $P_{m}$, which is computed using the cosine similarity of the Query-Key. Unlike L2P and Dual-Prompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathbf{A}$ to process the query features: $\alpha=\gamma\left(q(\mathbf{x})\odot\mathbf{A},\mathbf{K}\right)$, where $\mathbf{A}\subseteq\mathbb{R}^{D_{k}\times M}$ are learnable parameters. Finally, CODA-Prompt optimizes the prompt components $\mathbf{P}$, keys $\mathbf{K}$, attention $\mathbf{A}$ and the classifier. Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge. The final total loss function is:
$\min_{\mathbf{P},\mathbf{K},\mathbf{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathbf{P}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{K}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{A}\right)\right),$ (2)
where $\mathbf{P}$, $\mathbf{K}$, $\mathbf{A}$ refer to the prompt components and corresponding keys/attention vectors during tasks, respectively. The orthogonality penalty loss is $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix.
3.4 Prompt Fine-tuning
Based on the characteristics of large models, they possess outstanding feature extraction capabilities. However, due to the large model size, conventional fine-tuning approaches are not particularly effective. Current prompt-based continual learning methods employ a pre-trained Vision Transformer (ViT) as the backbone, with the backbone component frozen, which effectively preserves the internal hidden features of tasks. These methods adopt a prompt fine-tuning approach, shifting the training focus to the prompt pool. This approach only requires tuning a small portion of the model (the prompt embeddings) during the tasks, thereby effectively mitigating the issue of catastrophic forgetting. Based on the manner in which prompts are incorporated, prompt fine-tuning can be categorized into two types: prompt tuning and prefix tuning.
Figure 2: Our teacher-student model for continual distillation learning. The image is processed by the query function to generate a query, which is compared with $\mathbf{K}$ in the prompt pool. Based on different prompt extractor methods, the output prompts are divided into $\mathbf{P_{K}}$ and $\mathbf{P_{V}}$, which are inserted into the MSA layers across multiple blocks. The dashed boxes represent different types of optional distillation methods. Feature distillation uses the internal tokens output by intermediate blocks, and the class-token is used in the classification layer for logit distillation. Prompt Distillation adds an additional KD-Token and KD-Classifier to establish knowledge distillation with the teacher’s classification layer. In this case, the student model’s prediction is influenced by both the KD-Classifier and the Student-Classifier.
Prompt Tuning prepends the prompt to the input tokens. Equivalent to concatenating the prompt parameters to the input query ($\mathbf{h}_{Q}$), key ($\mathbf{h}_{K}$), and values ($\mathbf{h}_{V}$) of the MSA layer, this approach will increase the output length, where the output dimension is $\mathbb{R}^{(L+L_{p})\times D}$:
$f_{\text{Pro-T}}(P,\mathbf{h})=\text{MSA}([P;\mathbf{h}_{Q}],[P;\mathbf{h}_{K}],[P;\mathbf{h}_{V}]).$ (3)
Prefix Tuning divides the prompt $P$ into two parts, $P_{K}$, $P_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$, respectively:
$f_{\text{Pre-T}}(P,\mathbf{h})=\text{MSA}(\mathbf{h}_{Q},[P_{K};\mathbf{h}_{K}],[P_{V};\mathbf{h}_{V}]).$ (4)
This method keeps the output sequence length the same as the input. In this paper, since we need to insert prompts across multiple blocks, we employ the prefix method to avoid adding too many additional tokens. Then the prompts retrieved from the prompt pool are divided into two parts, $P_{K}$ and $P_{V}$.
4 Continual Knowledge Distillation
To study the CDL problem, we have designed a teacher-student model based on prompt-based continual learning methods, as illustrated in Fig. 2. Both the teacher model and the student model are some prompt-based continual learner such as L2P, DualPrompt or CODA-Prompt. Typically, the teacher’s backbone utilizes a larger and more accurate pretrained ViT model. The framework supports using different continual learners. We will study three different types of knowledge distillation within the framework: logit distillation, feature distillation and prompt distillation.
4.1 Prompt Extractor
To select a prompt, cosine similarity is used: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{m}\right)$ between the output of a query function $q(\mathbf{x})$ and a prompt key $\mathbf{k}_{m}$. Each key has a corresponding value in the prompt pool, forming key-value pairs. We can consider using different prompt extracting methods for different prompt-based methods. Then the final output prompts are divided into two parts, $P_{K}$ and $P_{V}$. Here, we use Prefix Tuning to insert $P_{K}$ and $P_{V}$ to the ViT backbone.
4.2 Logit Distillation
Logit distillation is one of the most classic forms of knowledge distillation. It aims to have a smaller student model learn the logits output of a larger or more accurate teacher model. In our CDL setup, we conducted experiments using two types of logit knowledge distillation methods: normal knowledge distillation (KD) [11] and Decoupled Knowledge Distillation (DKD) [39].
Normal Knowledge Distillation (KD): This approach was initially proposed by Hinton et al. [11], where distillation transfers the knowledge from a large, complex teacher model to a smaller student model, helping the latter to approximate the teacher model in terms of performance and accuracy. To achieve this, Hinton et al. designed a method where the logit output of the teacher model serves as “soft labels”, guiding the student model’s training. After passing through the softmax layer, the output values provide probabilities for each class.
$p_{i}=\frac{\exp(z_{i}/\tau)}{\sum_{j}\exp(z_{j}/\tau)},$ (5)
where $z_{i}$ is the logit, and $p_{i}$ is the predicted probability for class $i$. The soft target is the class probability distribution of teacher model. In the distillation process, a temperature parameter $\tau$, is introduced to smooth the output distribution of the model, making it easier for the student model to learn the subtle differences between classes.
In the teacher-student model shown in Fig. 2, the class tokens of the teacher model and the student model processed by the pre-trained ViT backbones are connected to the teacher classifier and student classifier to output their logits, respectively. The loss function of the student model for logit distillation is defined as:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{KD}}+\lambda\mathcal{L}_{\text{pool}},$ (6)
$\mathcal{L}_{\text{KD}}$ $=\tau^{2}\sum_{i}p_{i}^{\mathcal{T}}\log\left(\frac{p_{i}^{\mathcal{T}}}{p_{i}^{\mathcal{S}}}\right),$ (7)
where $\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)$ represents the cross-entropy classification loss used for learning with true label $y$, $g_{\phi}$ is the student classifier, and $f_{b}$ denotes the pre-trained ViT backbone and we only use the final class token into the classifier. $\mathcal{L}_{\text{KD}}$ represents the knowledge distillation loss, i.e., the KL divergence between the teacher’s probability distribution $p_{i}^{\mathcal{T}}$ and the student’s probability distribution $p_{i}^{\mathcal{S}}$. The $\mathcal{L}_{\text{pool}}$ is a loss function specific to the prompt pool. Different prompt-based continual learning methods have their respective prompt pool loss functions. For example, please refer to the loss functions in Equations (1) and (2) for L2P and CODA-Prompt. $\alpha$ and $\lambda$ are hyperparameters used to balance the weights of loss components.
Decoupled Knowledge Distillation (DKD) [39]: Unlike traditional knowledge distillation, which uses a unified KL divergence to measure the difference between the outputs of the student model and the teacher model, DKD separates the logit distillation loss into target class and non-target class components. DKD considers that the target and non-target classes contain different information and should be handled separately during training. By decoupling these components, DKD allows the student model to better capture the confidence on the target class while learning the distribution of the non-target classes. The student model loss function is formulated as follows:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)$
$+\alpha(\mathcal{L}_{\text{TCKD}}+\mathcal{L}_{\text{NCKD}})+\lambda\mathcal{L}_{\text{pool}},$ (8)
$\mathcal{L}_{\text{TCKD}}$ $=p_{t}^{\mathcal{T}}\log\left(\frac{p_{t}^{\mathcal{T}}}{p_{t}^{S}}\right)+p_{\setminus t}^{\mathcal{T}}\log\left(\frac{p_{t}^{\mathcal{T}}}{p_{\setminus t}^{\mathcal{S}}}\right)$ (9)
$\mathcal{L}_{\text{NCKD}}$ $=p_{\setminus t}^{\mathcal{T}}\sum_{\begin{subarray}{c}i\neq t\end{subarray}}\hat{p}_{i}^{\mathcal{T}}\log\left(\frac{\hat{p}_{i}^{\mathcal{T}}}{\hat{p}_{i}^{\mathcal{S}}}\right)$ (10)
where $\mathcal{L}_{\text{TCKD}}$ is the target class distillation loss, focused on aligning the student’s confidence with the teacher’s for the correct class. $\mathcal{L}_{\text{NCKD}}$ is the non-target class distillation loss, focused on matching the teacher and student model distributions for the incorrect classes. $[p_{t}^{\mathcal{T}},p_{\setminus t}^{\mathcal{T}}]$ represents the binary probabilities of the target class $p_{t}^{T}$ and all the other non-target classes $p_{\setminus t}^{\mathcal{T}}$ in teacher model, which can be calculated by Equation (5). $[p_{t}^{\mathcal{S}},p_{\setminus t}^{\mathcal{S}}]$ represents the binary probabilities in student model. Meanwhile, $\hat{p}_{i}^{\mathcal{T}}$ and $\hat{p}_{i}^{\mathcal{S}}$ are probability distributions among the non-target classes (without considering the $t$th class). Each element is calculated by:
$\hat{p}_{i}=\frac{\exp(z_{i}/\tau)}{\sum\limits_{j\neq t}\exp(z_{j}/\tau)}.$ (11)
4.3 Feature Distillation
Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model. It can leverage richer, layer-wise information within the teacher model to guide the student model’s learning. The student model can benefit from an understanding of feature relationships. In the CDL model in Fig. 2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features. This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model. It is worth noting that during knowledge distillation, the backbone of the student model remains frozen while processing the internal tokens, which is a characteristic of prompt-based continual learning methods. The frozen part is crucial for mitigating catastrophic forgetting during the continual learning process. In this paper, we study two methods based on the handling of internal tokens: FitNets [26] and Review Knowledge Distillation (ReviewKD) [4].
FitNets [26]: This method is one of the most classic feature distillation methods, aimed at adding a distillation loss to the intermediate layers. It uses the intermediate feature representations of the teacher model as hints to guide the student model’s learning. Here, we select the output of the last block of the teacher model as the hint. Similarly, the student model selects the output of the corresponding block for learning, constructing the feature distillation loss. Finally, the total loss function for the student model is
$\mathcal{L}_{\text{S}}$ $=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{hint}}+\lambda\mathcal{L}_{\text{pool}},$ (12)
$\mathcal{L}_{\text{hint}}$ $=\left\|f_{b-1}^{\mathcal{T}}(\mathbf{x})-F_{M}(f_{b-1}^{\mathcal{S}}(\mathbf{x}))\right\|^{2},$ (13)
where $\mathcal{L}_{\text{hint}}$ is the feature distillation loss, calculated using the Mean Squared Error (MSE). $\alpha$ and $\lambda$ are used to balance the weights of loss components. $f_{b-1}$ indicates the feature output (internal tokens) after the last block of the ViT model. The student feature is transformed into the same size as the teacher feature with the mapping layer $F_{M}$, which is simply a fully-connected layer in the network.
ReviewKD [4]: It innovates by “reviewing” multiple hidden layers from both the teacher and student models, offering a more comprehensive approach to capturing hierarchical features across the entire model. It reviews the multiple layers utilizing the concept of residual learning. For instance, the feature from $n_{th}$ block of the student is aggregated with the feature from $(n-1)_{th}$ block of the student to mimic the feature from $(n-1)_{th}$ block of the teacher. The total loss function of the student in ReviewKD is
$\mathcal{L}_{\text{S}}=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{RKD}}+\lambda\mathcal{L}_{\text{pool}},$ (14)
$\mathcal{L}_{\text{RKD}}=\mathcal{D}(\mathbf{F}^{\mathcal{S}}_{n},\mathbf{F}^{\mathcal{T}}_{n})+\sum_{j=n-1}^{1}\mathcal{D}\left(\mathcal{U}(\mathbf{F}_{j}^{\mathcal{S}},\mathbf{F}_{j+1}^{\mathcal{S}}),\mathbf{F}^{\mathcal{T}}_{j}\right),$ (15)
where $\mathcal{L}_{\text{RKD}}$ is the reviewKD loss, and $\mathcal{D}$ is L2 distance between the student features and teacher features. All student features in the equations have passed through the mapping layer $F_{M}$ to make them the same dimension as the teacher features. $\mathbf{F}_{j}^{\mathcal{S}}$ and $\mathbf{F}_{j}^{\mathcal{T}}$ are the features output by the student model and teacher model, respectively, after passing through $j$ blocks. $\mathbf{F}_{j+1}^{\mathcal{S}}$ represents the fused student features at the $(j+1)_{th}$ block. $\mathcal{U}$ is a module used to fuse features, which performs a weighted combination of the two input features. $n$ is the total number of blocks in the ViT backbone. Therefore, in the student model, the fused features obtained at each block are passed to the next higher block to form new feature fusion.
4.4 Prompt Distillation
In the traditional knowledge distillation, the backbone of the student model is trainable, allowing it to learn and approach the teacher model. However, in the prompt-based methods, the backbone is frozen, which limits the update of intermediate features extracted by the student model, making it challenging to effectively learn knowledge from the teacher model’s features. Similarly, the logits are also impacted during training updates.
In continual learning, as tasks increase, the model seeks to retain information from previous tasks. Prompt-based continual learning leverages a pretrained and frozen ViT model to preserve knowledge across tasks. However, in knowledge distillation, the student model aims to be fully trained and adjusted to approach the teacher model. These two objectives are conflicted with each other, which creates mutual interference between the continual learning process and the knowledge distillation process. Subsequently, it affects the performance of the student model. Here, we introduce a new concept Continual Distillation Interference (CDI) to describe this phenomenon. CDI is commonly observed in various continual learning methods that rely on large pretrained models as we will show in our experiments.
To address the continual interference problem of distillation, we explored a prompt distillation method as motivated by DeiT [29]. As shown in Fig. 2, we introduce a learnable prompt called KD-Token in the student model, which operates alongside the class token to assist in classification. The KD-Token in the student model can be considered as learning knowledge from the class token in the teacher model. After passing through the ViT backbone, the KD token is input into a newly added KD classifier. The KD classifier has two roles. First, it is used for knowledge distillation with the teacher model. Second, the output of the KD classifier and the output of the student classifier are averaged to generate the final classification result.
Therefore, the KD-Token is specifically designed to assist in knowledge distillation, while the original prompt tokens $P_{K}$ and $P_{V}$, obtained from the prompt pool, are used to facilitate the continual learning process, each serving its respective role. The purpose of this method is to mitigate the CDI problem in prompt-based continual learning. The training loss function of the student model is the same as the one in Equation (6). However, the KD loss is between the KD classifier and teacher classifier. Here, we highlight the difference by:
$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}^{\mathcal{S}}(f_{b}(\mathbf{x})),y)$
$+\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(\mathbf{F}^{\mathcal{S}}_{n}[kd]),g_{\phi}^{\mathcal{T}}(\mathbf{F}^{\mathcal{T}}_{n}[cls]))+\lambda\mathcal{L}_{\text{pool}},$ (16)
where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively. $g_{\phi}^{\mathcal{T}}$ is the teacher classifier. $n$ is the total number of blocks in the ViT backbone. $\mathbf{F}^{\mathcal{S}}_{n}[kd]$ represent the KD-Token after the $n_{\text{th}}$ block of the student model. $\mathbf{F}^{\mathcal{T}}_{n}[cls]$ denotes the class token in the teacher model. The knowledge distillation loss $\mathcal{L}_{\text{KD}}$ (Equation (7)) uses KL divergence between the KD classifier and teacher classifier. $\mathcal{L}_{\text{pool}}$ represents the loss function on the prompt pool across different continual learning methods, and $\alpha$ and $\lambda$ are balancing weights.
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
ViT-Base ViT-Small KD [11] $69.23\pm 0.05$ $8.57\pm 0.39$ $82.39\pm 0.25$ $7.33\pm 0.21$
ViT-Base ViT-Small DKD [39] $68.92\pm 0.07$ $8.39\pm 0.36$ $82.27\pm 0.20$ $7.81\pm 0.14$
ViT-Base ViT-Small FitNets [26] $69.87\pm 0.04$ $7.38\pm 0.36$ $81.83\pm 0.05$ $8.83\pm 0.48$
ViT-Base ViT-Small ReviewKD [4] $70.19\pm 0.16$ $7.68\pm 0.01$ $82.20\pm 0.41$ $7.54\pm 0.03$
ViT-Base ViT-Small KD-Token [29] $\mathbf{70.94\pm 0.33}$ $\mathbf{6.17\pm 0.07}$ $\mathbf{83.72\pm 0.08}$ $\mathbf{6.40\pm 0.06}$
$\varnothing$ ViT-Base $75.78\pm 0.64$ $5.70\pm 0.66$ $86.16\pm 0.17$ $5.63\pm 0.25$
ViT-Large ViT-Base KD [11] $\mathbf{77.51\pm 0.01}$ $\mathbf{3.44\pm 0.14}$ $87.41\pm 0.05$ $5.45\pm 0.06$
ViT-Large ViT-Base DKD [39] $76.70\pm 0.17$ $4.84\pm 0.12$ $85.42\pm 0.31$ $6.55\pm 0.16$
ViT-Large ViT-Base FitNets [26] $74.55\pm 0.14$ $6.81\pm 0.15$ $85.95\pm 0.25$ $6.56\pm 0.02$
ViT-Large ViT-Base ReviewKD [4] $75.72\pm 0.27$ $4.14\pm 0.05$ $86.21\pm 0.61$ $5.64\pm 0.40$
ViT-Large ViT-Base KD-Token [29] $77.28\pm 0.29$ $4.63\pm 0.16$ $\mathbf{88.10\pm 0.33}$ $\mathbf{5.39\pm 0.19}$
Table 1: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on CODA-Prompt [28].
5 Experimental Results
5.1 Experimental Benchmarks and Settings
We utilize the CIFAR-100 [16] and ImageNet-R [10] datasets in a class-incremental continual learning setting for our experiments. Following previous works, we divide the ImageNet-R and CIFAR-100 into 10 tasks, where each task contains 10 classes. We experiment with three prompt-based continual learning models, i.e., L2P [34], DualPrompt [33] and CODAPrompt [28]. In DualPrompt, the G-prompts are set in the first two layers with a prompt length of 5, and the E-prompts are set in layers 3 to 5. CODA-Prompt sets the prompt length to 8, with a total of 100 prompt components inserted into the first 5 layers. This setup is the same for all ViT models. Our experiments tested two teacher-student knowledge distillation pairs. One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6]. For different knowledge distillation methods, the hyperparameter $\alpha$ for the distillation loss is set to 0.5.
We use two metrics to evaluate the models: accuracy and forgetting rate [21]. The accuracy refers to the average accuracy of all tasks after completing all 10 tasks. The forgetting rate reflects the influence of learning a new task on previously completed tasks. A larger forgetting signifies a more negative impact of the continual learning model. Please see the supplementary material for the details.
5.2 Continual Distillation Results
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small + L2P [34] $61.95\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
$\varnothing$ ViT-Small + DualPrompt [33] $64.27\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
$\varnothing$ ViT-Small + CODA-Prompt [28] $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
ViT-Base ViT-Small KD-Token [29] + L2P [34] $65.09\pm 1.05$ $5.37\pm 0.13$ $79.78\pm 0.14$ $7.07\pm 0.52$
ViT-Base ViT-Small KD-Token [29] + DualPrompt [33] $66.81\pm 0.42$ $\mathbf{4.55\pm 0.31}$ $81.32\pm 0.36$ $\mathbf{5.67\pm 0.59}$
ViT-Base ViT-Small KD-Token [29] + CODA-Prompt [28] $\mathbf{70.94\pm 0.33}$ $6.17\pm 0.07$ $\mathbf{83.72\pm 0.08}$ $6.40\pm 0.06$
$\varnothing$ ViT-Base + L2P [34] $71.59\pm 1.20$ $5.65\pm 0.69$ $83.02\pm 0.47$ $6.06\pm 0.47$
$\varnothing$ ViT-Base + DualPrompt [33] $72.44\pm 1.00$ $3.80\pm 1.23$ $84.66\pm 0.87$ $5.91\pm 0.34$
$\varnothing$ ViT-Base + CODA-Prompt [28] $75.78\pm 0.64$ $5.70\pm 0.66$ $86.16\pm 0.17$ $5.63\pm 0.25$
ViT-Large ViT-Base KD-Token [29] + L2P [34] $71.89\pm 0.03$ $\mathbf{3.42\pm 0.36}$ $85.2\pm 0.28$ $\mathbf{4.58\pm 0.55}$
ViT-Large ViT-Base KD-Token [29] + DualPrompt [33] $73.63\pm 0.26$ $3.5\pm 0.25$ $85.32\pm 0.02$ $5.64\pm 0.51$
ViT-Large ViT-Base KD-Token [29] + CODA-Prompt [28] $\mathbf{77.28\pm 0.29}$ $4.63\pm 0.16$ $\mathbf{88.10\pm 0.33}$ $5.39\pm 0.19$
Table 2: Results of prompt distillation using KD token on the ImageNet-R dataset and the CIFAR-100 dataset with different teacher-student models and different continual learning models.
Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner. The complete results for the L2P and DualPrompt methods are provided in the supplementary material. When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation. The student model is trained with the continual learning task only. We hope that the knowledge distillation methods can improve over this baseline by increasing accuracy and reducing forgetting rate.
From the table, we observe that, prompt distillation using KD-Token achieves the best overall performance on the two datasets. The other knowledge distillation methods exhibit certain limitations to varying extents. For example, KD with ViT-Small leads to increased forgetting, and feature distillation methods such as FitNets and ReviewKD even reduce accuracy in some cases. Although KD, DKD, FitNets, and ReviewKD may improve continual learning performance in some situations, their effects are not significant. They cannot simultaneously ensure both accuracy and reduced forgetting. This exemplifies the continual distillation interference problem mentioned in Section  4.4.
In contrast, the prompt distillation method using KD-Token demonstrates a more significant improvement for the student model. Table 2 presents the distillation results of the prompt distillation (KD-Token) method combined with L2P, DualPrompt and CODA-Prompt. We observe that, in all cases, the method with KD-Token distillation outperforms the original model, as reflected in the accuracy and forgetting rate metrics, validating the effectiveness of this approach. In addition, among the three continual learning models, CODA-Prompt achieves the best performance with or without knowledge distillation.
Balance $\mathbf{\alpha}$ 00 $0.1$ $0.3$ $0.5$ $0.7$ $0.9$ $1.0$
Avg. Acc ($\uparrow$) $67.44$ $68.96$ $70.06$ $\mathbf{70.94}$ $69.34$ $68.24$ $68.06$
Forgetting ($\downarrow$) $8.52$ $7.11$ $6.46$ $\mathbf{6.17}$ $6.22$ $6.40$ $6.85$
Table 3: Ablation study of the parameter $\alpha$ based on CODA-Prompt [28] with the Vit-Base to ViT-Small teacher-student model for prompt distillation using a KD-Token.
5.3 Continual Distillation Interference
The prompt distillation method shifts the distillation training process to the KD-Token and KD classifier, thereby enhancing the efficiency of the distillation. Based on our analysis, the freezing of the backbone results in suboptimal training outcomes for logit distillation and feature distillation. However, making the backbone trainable can result in a severe forgetting problem.
To investigate the impact of adding trainable components to the backbone, we conducted a comparative experiment involving the following three models: 1) Normal KD with an additional trainable MLP layer added after the last block of the backbone (Add-MLP); 2) Prompt distillation wiht a KD-Token (Add-KD-Token); and 3) the original model without distillation (No KD). We distill from ViT-Base to ViT-Small using CODA-Prompt. After training on all tasks, we recorded the test accuracy for each task. The results are shown in Fig. 3. It can be observed that KD with an MLP improves the distillation effectiveness for the current task and the most recent task, but results in severe forgetting for earlier tasks. Essentially, the objective of continual learning is to retain information from previous tasks as much as possible (model rigidity), while knowledge distillation aims for the student model to learn from the teacher model as effectively as possible (model adaptability). This introduces a mutual interference between continual learning and distillation, which we refer to as Continual Distillation Interference (CDI). The performance of Add_MLP demonstrates the existence of this problem. The KD-Token method attempts to address this problem by adding a KD-Token, shifting the distillation task to this prompt. We observe that the KD-Token method not only improves the accuracy of the current task through knowledge distillation but also maintains improved performance on previous tasks.
5.4 Ablation Studies
[…]
Figure 3: After completing all tasks, the test accuracy results for each task are recorded. The CDL method uses ViT-Base to ViT-Small, with CODA-Prompt as the baseline.
6 Conclusion and Discussion
We introduce the problem of Continual Distillation Learning (CDL), which considers knowledge distillation in the continual learning setup. We emperically studied three types of knowledge distillation methods in CDL, i.e., logit distillation, feature distillation, and prompt distillation, where three different prompt-based continual learning methods (L2P, DualPrompt, and CODA-Prompt) are utilized for this study. We identified the presence of the Continual Distillation Interference (CDI) phenomenon in the process of CDL. The experimental results demonstrate that the KD-Token method outperforms the other distillation methods, consistently helping the student model improve performance in different scenarios. The primary purpose of this paper is to serve as an inspiration in the hope that more research on CDL will emerge in the future.
Limitations. Knowledge distillation relies on a well-performing teacher model. Prompt-based continual learning methods can obtain teacher models by using large ViT backbones. However, for traditional continual learning methods that rely on CNN backbones (e.g., EWC [15] and iCaRL [24]), it is not straightforward to obtain a well-performing teacher model. This is why we limit our study to prompt-based continual learning. In addition, since each task requires training a larger teacher model, the total training time increases.
Acknowledgments
This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, and the National Science Foundation (NSF) under Grant No. 2346528.
Supplementary Material
{strip}
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $61.95\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
ViT-Base ViT-Small KD [11] $63.36\pm 0.04$ $6.74\pm 0.10$ $77.92\pm 0.23$ $8.30\pm 0.29$
ViT-Base ViT-Small DKD [39] $62.91\pm 0.27$ $6.55\pm 0.17$ $78.21\pm 0.12$ $9.13\pm 0.07$
ViT-Base ViT-Small FitNets [26] $64.29\pm 0.09$ $6.37\pm 0.17$ $79.56\pm 0.39$ $\mathbf{5.89\pm 0.36}$
ViT-Base ViT-Small ReviewKD [4] $63.64\pm 0.34$ $6.36\pm 0.58$ $78.50\pm 0.39$ $8.04\pm 0.75$
ViT-Base ViT-Small KD-Token [29] $\mathbf{65.09\pm 1.05}$ $\mathbf{5.37\pm 0.13}$ $\mathbf{79.78\pm 0.14}$ $7.07\pm 0.52$
$\varnothing$ ViT-Base $71.59\pm 1.20$ $5.65\pm 0.69$ $83.02\pm 0.47$ $6.06\pm 0.47$
ViT-Large ViT-Base KD [11] $74.12\pm 0.42$ $4.60\pm 0.55$ $80.96\pm 0.34$ $10.17\pm 0.67$
ViT-Large ViT-Base DKD [39] $\mathbf{74.58\pm 0.01}$ $4.69\pm 0.06$ $83.29\pm 0.24$ $4.99\pm 0.17$
ViT-Large ViT-Base FitNets [26] $70.39\pm 0.23$ $5.84\pm 0.06$ $83.60\pm 0.02$ $5.21\pm 0.71$
ViT-Large ViT-Base ReviewKD [4] $72.17\pm 0.26$ $6.11\pm 0.08$ $83.12\pm 0.65$ $7.97\pm 0.27$
ViT-Large ViT-Base KD-Token [29] $71.89\pm 0.03$ $\mathbf{3.42\pm 0.36}$ $\mathbf{85.20\pm 0.28}$ $\mathbf{4.58\pm 0.55}$
Table 4: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on L2P [34].
7 Implementation Details
Our experiments use two metrics to evaluate the models: accuracy and forgetting rate. Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Equation (17), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$. $T$ is the total number of tasks. The forgetting rate (Equation (18)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks. A higher value signifies a more negative impact of the continual learning model.
$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i}$ (17)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}$ (18)
For all experiments, we utilized the Adam optimizer [14]. The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs. The learning rate was set to $l=0.001$. In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$. Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
8 Continual Distillation Results
Tables Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning and 5 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [34] and DualPrompt [33] as the continual learners. The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners. Notably, it achieves the highest accuracy in the L2P and DualPrompt models on the CIFAR-10 dataset. Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.
Methods Split ImageNet-R Split CIFAR-100
Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
$\varnothing$ ViT-Small $64.27\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
ViT-Base ViT-Small KD [11] $65.68\pm 0.06$ $7.26\pm 0.29$ $80.16\pm 0.54$ $5.76\pm 0.18$
ViT-Base ViT-Small DKD [39] $65.44\pm 0.05$ $7.27\pm 0.23$ $80.44\pm 0.46$ $6.96\pm 0.44$
ViT-Base ViT-Small FitNets [26] $66.20\pm 0.14$ $5.93\pm 0.17$ $80.70\pm 0.17$ $5.73\pm 0.21$
ViT-Base ViT-Small ReviewKD [4] $65.69\pm 0.91$ $6.56\pm 0.46$ $80.33\pm 0.23$ $5.86\pm 0.53$
ViT-Base ViT-Small KD-Token [29] $\mathbf{66.81\pm 0.42}$ $\mathbf{4.55\pm 0.31}$ $\mathbf{81.32\pm 0.36}$ $\mathbf{5.67\pm 0.59}$
$\varnothing$ ViT-Base $72.44\pm 1.00$ $3.80\pm 1.23$ $84.66\pm 0.87$ $5.91\pm 0.34$
ViT-Large ViT-Base KD [11] $73.90\pm 0.14$ $\mathbf{3.31\pm 0.04}$ $83.4\pm 0.74$ $8.23\pm 1.13$
ViT-Large ViT-Base DKD [39] $\mathbf{75.24\pm 0.33}$ $4.15\pm 0.23$ $84.93\pm 0.16$ $\mathbf{4.95\pm 0.12}$
ViT-Large ViT-Base FitNets [26] $71.23\pm 0.04$ $5.71\pm 0.55$ $83.12\pm 0.86$ $8.33\pm 1.36$
ViT-Large ViT-Base ReviewKD [4] $72.19\pm 0.01$ $5.72\pm 0.20$ $84.11\pm 0.92$ $5.19\pm 0.33$
ViT-Large ViT-Base KD-Token [29] $73.63\pm 0.26$ $3.50\pm 0.25$ $\mathbf{85.32\pm 0.02}$ $5.64\pm 0.51$
Table 5: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on DualPrompt [33].

### v3

@@
-Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning
+Continual Distillation Learning: Knowledge Distillation in Prompt-based Continual Learning
 Abstract
-Knowledge Distillation (KD) focuses on using a teacher model to improve a student model.
-Traditionally, KD is studied in an offline fashion, where a training dataset is available before learning.
-In this work, we introduce the problem of Continual Distillation Learning (CDL) that considers KD in the Continual Learning (CL) setup.
-A teacher model and a student model need to learn a sequence of tasks, and the knowledge of the teacher model will be distilled to the student to improve the student model in an online fashion.
-The CDL problem is valuable to study since for prompt-based continual learning methods, using a larger vision transformer (ViT) leads to better performance in continual learning.
-Distilling the knowledge from a large ViT to a small ViT can improve inference efficiency for prompt-based CL models.
-To this end, we conducted experiments to study the CDL problem with three prompt-based CL models, i.e., L2P, DualPrompt and CODA-Prompt, where we utilized logit distillation, feature distillation and prompt distillation for knowledge distillation from a teacher model to a student model.
-Our findings of this study can serve as baselines for future CDL work.
+We introduce the problem of continual distillation learning (CDL) in order to use knowledge distillation (KD) to improve prompt-based continual learning (CL) models.
+The CDL problem is valuable to study since the use of a larger vision transformer (ViT) leads to better performance in prompt-based continual learning.
+The distillation of knowledge from a large ViT to a small ViT can improve the inference efficiency for prompt-based CL models.
+We empirically found that existing KD methods such as logit distillation and feature distillation cannot effectively improve the student model in the CDL setup.
+To this end, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model.
+We demonstrate that our KDP method effectively enhances the distillation performance in comparison to existing KD methods in the CDL setup.
 [^1] Project website https://irvlutd.github.io/CDL/
@@
 Continual Learning (CL) [32] designs models that can continuously learn new tasks without forgetting previously learned tasks.
-In class-incremental continual learning, data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training.
+For example, in class-incremental continual learning [26], the data of new classes arrive sequentially, and the model needs to learn to recognize these classes sequentially during training.
 In testing, the model will be tested on all the seen classes.
 Therefore, a good CL model should learn new classes without forgetting.
-Traditionally, the class-incremental setting is tackled using a rehearsal buffer [24, 17, 21, 3], which stores training examples from previous classes.
-During training, data from new tasks and data from the rehearsal buffer are used to update the model weights.
-With the help of the rehearsal buffer, the model can continually refresh its memory of previous classes.
-However, due to the size limit of the rehearsal buffer, these models still suffer from the catastrophic forgetting problem [22].
-Recently, a new paradigm for CL based on learning for prompt receives more attentions [34, 33, 28, 31].
-Instead of using rehearsal buffers, these methods focus on learning prompts for pre-trained large models such as vision transformers for continual learning.
-For example, L2P [34] learns a prompt pool for a set of tasks using pre-trained ViTs [6], where different tasks learn to use different prompts in the prompt pool.
-DualPrompt [33] learns two types of prompts, i.e., a global prompt for all tasks and task-specific prompts.
-By leveraging the pre-trained ViT models, these models achieve better performance compared to traditional rehearsal buffer-based methods.
+With recent advances in vision models, vision transformers (ViTs) [6] have demonstrated their advantages over convolutional neural networks (CNNs).
+Similarly, continual learning research has evolved from traditional CNN-based methods (e.g., using ResNet [11]) to the latest prompt-based CL methods that leverage ViTs as the backbone, which now represent the state-of-the-art CL models.
+Unlike traditional CL models where the backbone is trained, prompt-based CL methods such as L2P [35], DualPrompt [34] and CODA-Prompt [30], utilize a pre-trained ViT as the backbone while keeping it frozen during training.
+Instead, they optimize a prompt pool designed for continual learning.
+Using pre-trained ViT models and shifting learning to prompts, these methods achieve better performance compared to CNN-based continual learning approaches such as iCaRL [26] and LWF [19].
+Figure 1: Comparison between two knowledge distillation scenarios. (a) Traditional KD with an offline dataset. (b) The CDL problem introduced in this work.
 We noticed that for prompt-based CL methods using ViTs, larger ViTs [6] achieve better performance.
@@
 However, larger models introduce more computation during inference.
-To address this problem, we study the problem of Knowledge Distillation (KD) from large to small models in the continual learning setting.
-We name this problem Continual Distillation Learning (CDL).
-Figure 1: Comparison between three different knowledge distillation scenarios.
-(a) Traditional KD with an offline dataset.
-(b) KD in rehearsal-based CL models.
-(c) The CDL problem introduced in this work.
+To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in prompt-based continual learning.
+We name this problem Continual Distillation Learning (CDL), which aims to improve the performance of small models by using large models as teachers in continual learning.
+We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as the iCaRL [26] model or the LWF [19] model (see Sec.
+7 in supplementary material).
+These CNN-based models update the backbones during training.
+Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem in continual learning.
 The CDL problem is different from the previous knowledge distillation setup.
-We illustrate these differences in Fig.
+We illustrate the differences in Fig.
 1.
-First, in the traditional KD setup, an offline dataset is used for knowledge distillation (Fig.
+In the traditional KD setup [13], an offline dataset is used for knowledge distillation (Fig.
 1(a)).
@@
 The setup is not the continual learning setup.
-Second, KD is also used in the rehearsal-based CL methods such as iCaRL [24] (Fig.
-1(b)).
-These methods maintain a rehearsal buffer of previous data.
-When updating the model parameters for the current task, KD ensures that the prediction on the previous data should be similar to the prediction of the old model.
-In this case, the old model before updating is the teacher model.
-Lastly, the CDL problem we study in this work is illustrated in Fig.
-1(c).
+The CDL problem that we study in this work is illustrated in Fig.
+1(b).
 Given a new task in continual learning, a teacher model is first updated.
 Then a student model is updated based on the data of the new task and the teacher model.
-Consequently, a better student model can be trained by knowledge distillation from the teacher model.
-To study the CDL problem, we have designed a prompt-based continual learning method that consists of a teacher model and a student model.
-In our method, we jointly learn the teacher model and the student model.
-Both models are prompt-based CL models, where the backbones are frozen and the prompts are learnable.
-When training data for new tasks come, the teacher model is first updated as other prompt-based CL models in the literature, and then the student model is updated.
-Our method is flexible to be used with different prompt-based CL models and different mechanisms for knowledge distillation.
-Therefore, we conducted an empirical study of the CDL problem using three prompt-based CL models, i.e., L2P [34], DualPrompt [33] and CODA-Prompt [28].
-For each model, we considered different teacher-student configurations.
-For example, using ViT large as the teacher model and ViT base as the student model.
-In addition, we studied the effect of three types of knowledge distillation mechanism in the CDL setup, i.e.
-logit distillation, feature distillation, and prompt distillation.
-We found that prompt distillation works the best, due to the prompt learning nature of these models.
+Consequently, a better student model can be trained by distilling knowledge from the teacher model.
+In general, this is a task-incremental learning process, where both the teacher and student models can only access the data for the current task at each step.
+To explore the new CDL problem, we first conducted an empirical study by applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4] and distillation token [31], to three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34], and CODA-Prompt [30].
+For each model, we consider different teacher-student configurations.
+For example, we can use ViT-Large as the teacher model and ViT-Base as the student model.
+Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool, where the selection of prompts relies on a key-query mechanism.
+However, for knowledge distillation, this mechanism is not effective.
+Since each task independently reselects prompt components from the prompt pool, the knowledge embedded within them from the teacher model cannot be transferred to the next task.
+Our experimental results also show that the performance improvements achieved by these knowledge distillation methods are not significant.
+Previous KD methods are not effective in the CDL setting.
+To address this limitation, we propose a novel method named Knowledge Distillation based on Prompts (KDP).
+KDP inserts globally accessible prompts, specifically designed for knowledge distillation, into the frozen ViT backbone of the student model.
+We name them KD prompts.
+These prompts are independent of the key-query mechanism of the prompt pool.
+They are not limited to specific tasks and can facilitate cross-task distillation.
+They serve as auxiliary information to guide the learning process of the student model.
+Experimental results demonstrate that this novel approach effectively addresses the CDL problem and enhances distillation performance compared to previous KD methods.
+Furthermore, we conducted an ablation study on the total number of KD prompts inserted into the backbone.
+The results show that the introduction of KD prompts into each ViT block significantly improves the alignment of the feature structures between layers of the teacher model and the student model, leading to better performance of the student model.
 In summary, our contributions are as follows.
-  • We introduce the problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
-  • We designed a method for CDL that can be used with different prompt-based CL models and different knowledge distillation mechanism.
-  • We conducted extensive experiments and analyses of three prompt-based CL models with three types of KD mechanism on two commonly-used benchmarks for continual learning.
+  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
+  • We conducted an empirical study by applying different knowledge distillation methods to three prompt-based CL approaches and identify the limitations of these distillation methods in the CDL setting.
+  • To address the unique challenges of the CDL problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches.
 2 Related Work
-Continual Learning.
+2.1 Continual Learning
 The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting [24] in the case of incremental tasks.
@@
 Architecture-based approaches [38, 15, 36, 7, 25] mainly focus on constructing a special model for continual learning.
-Replay-based approaches [24, 17, 12, 21, 1, 3, 2] use the memory buffer to store and replay old task data in learning the current task.
+Replay-based approaches [26, 19, 14, 23, 1, 3, 2] use a memory buffer to store and replay old task data in learning the current task.
 Some recent works aim to address the issue of catastrophic forgetting without relying on rehearsal memory, which are referred as rehearsal-free methods [10, 8, 5].
-The aforementioned methods do not leverage pre-trained networks.
-They usually learn the tasks from scratch.
-Prompt-based Continual Learning.
+The above methods are primarily based on CNN backbones, and a larger backbone model may not achieve better performance.
+2.2 Prompt-based Continual Learning
 Inspired by the use of prompts in natural language processing [21], prompt-based continual learning methods employ large vision transformer (ViT) models [6].
-These methods freeze the pre-trained backbone and shift learning to the prompts of the ViTs.
-For example, L2P [34] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool to select the corresponding prompts.
+These methods freeze the pre-trained backbone and shift learning to prompts of the ViTs.
+For example, L2P [35] learns a prompt pool of key-prompt pairs, then selects the optimal prompt in the pool for a given input by matching the input with the keys in the pool.
 Based on this idea, DualPrompt [34] introduces the concept of using a general prompt and an expert prompt.
@@
 It uses a weighted sum of the prompt components to obtain the final prompt.
+CPrompt [9] introduces two main components during training: Classifier Consistency Learning (CCL) and Prompt Consistency Learning (PCL).
 In prompt-based methods, the size of the backbone matters.
 Larger backbone models achieve better performance.
-This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to speed up inference time.
-Knowledge Distillation (KD).
-The purpose of knowledge distillation is to leverage large models to assist smaller models in fully exploiting their potential.
-The larger model, termed the teacher, aids in training the smaller model, referred to as the student.
-According to the distillation location, traditional methods can be divided into logit distillation [11, 39] and feature distillation [18, 26, 4].
-With the recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [29, 36].
+This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to improve small models with fast inference time.
+2.3 Knowledge Distillation
+The purpose of knowledge distillation is to leverage larger models to assist smaller models to fully exploit their potential.
+The larger model, termed the teacher, aids in training the smaller model, termed the student.
+Traditional KD methods can be classified into logit distillation [13, 40] and feature distillation [20, 28, 4].
+With recent rapid advancements in large models, new research has directed the focus of knowledge distillation towards vision transformers [31, 37].
 However, KD for prompt-based continual learning models has not yet been explored, which is the focus of this work.
 3 Prerequisites
-3.1 L2P
-L2P [34] and DualPrompt [33] use a pre-trained ViT model as the backbone and keep the backbone frozen.
-Both methods use a prompt pool to encode information about tasks: $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}$, where $P_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompts.
-In order to select prompts for different tasks, each prompt is associated with a learnable key $\{(\mathbf{k}_{1},P_{1}),(\mathbf{k}_{2},P_{2}),\ldots,(\mathbf{k}_{M},P_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ and $\mathbf{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys.
-Given an input image $\mathbf{x}$, a query function $q(\cdot)$ is used to encode the input image.
-The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(x),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function.
-Then the top $k$ prompts from the prompt pool $\mathbf{P}$ based on the establishment of key-value pairs are selected.
-These prompts are used in the ViT model to classify the input.
-In this way, given an input-label pair $(\mathbf{x},y)$, L2P optimizes the trainable prompt keys and values, and the final classification layer.
-The training loss function is defined as
-$\min_{\mathbf{P},\mathbf{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (1)
-where $g_{\phi}$ is the classifier, $f_{b}$ is the pre-trained ViT backbone, $\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification.
+3.1 Continual Learning Setting
+In continual learning, a model is required to learn a sequence of tasks, where the data from these tasks arrive on time.
+We denote a sequence of tasks as $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{T}\}$, where $T$ is the number of tasks.
+The $t$th task $\mathcal{D}_{t}=\{(\mathbf{x}_{i}^{t},y_{i}^{t})\}_{i=1}^{n_{t}}$ consists of pairs of input sample $\mathbf{x}_{i}^{t}\in\mathcal{X}$ and its label $y_{i}^{t}\in\mathcal{Y}$, where $n^{t}$ is the number of samples for the $t$th task.
+In this work, we consider class-incremental learning, where each task consists of a fixed number of non-overlapping classes.
+In training, a model learns these tasks one by one.
+Data from the previous tasks are not available anymore when training future tasks.
+In testing, the model is evaluated by testing samples from all classes.
+Figure 2: Our teacher-student model for continual distillation learning.
+The green and red dashed boxes represent different types of optional distillation methods.
+The blue dashed box represents our KDP method, where the components inside are the trainable modules to be added.
+3.2 Prompt-based Methods
+Prompt-based continual learning methods use pre-trained ViT backbones and freeze the backbones during training. Learning is shifted to trainable prompts.
+L2P: L2P [35] uses a prompt pool to encode information about tasks: $\mathcal{P}=\{\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}\}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L_{p}\times D}$ with token length $L_{p}$ and embedding size $D$, and $M$ is the total number of prompt components.
+To select prompts for different tasks, each prompt component is associated with a learnable key $\{(\mathbf{k}_{1},\mathbf{P}_{1}),(\mathbf{k}_{2},\mathbf{P}_{2}),\ldots,(\mathbf{k}_{M},\mathbf{P}_{M})\}$, where $\mathbf{k}_{i}\in\mathbb{R}^{D_{k}}$ with embedding size $D_{k}$ and $\mathcal{K}=\{\mathbf{k}_{i}\}_{i=1}^{M}$ denotes all the keys.
+Given an input image $\mathbf{x}\in\mathcal{X}$, a query function $q(\cdot)$ is used to encode the input image.
+The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function.
+Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected:
+$\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}),$ (1)
+where TopK indicates the selection function based on cosine similarity.
+$\mathbf{P}_{\mathbf{x}}\in\mathbb{R}^{K\times L_{p}\times D}$ is the final set of prompt components extracted from the prompt pool to assist continual learning.
+The training loss function for a pair $(\mathbf{x},y)$ is defined as:
+$\min_{\mathcal{P},\mathcal{K},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\lambda\sum_{\mathbf{k}_{i}\in\mathbf{K}_{\mathbf{x}}}\gamma(q(\mathbf{x}),\mathbf{k}_{i}),$ (2)
+where $g_{\phi}$ is the classifier with parameter $\phi$.
+$f_{b}$ is the pre-trained ViT backbone, which includes the selected prompt components $\mathbf{P}_{\mathbf{x}}$ for continual learning.
+$\mathbf{K}_{\mathbf{x}}$ denotes the selected keys for input $\mathbf{x}$, and $\mathcal{L}$ denotes the softmax cross-entropy loss for classification.
 The final class token of the ViT is used for the classifier.
-Therefore, the first part in Eq.
-(1) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathbf{P}$, key set $\mathbf{K}$ and the classifier parameter $\phi$.
-The second term minimizes the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
-3.2 DualPrompt
-The DualPrompt method [33] supplements insertable prompts based on the L2P method.
+Therefore, the first term in Eq.
+(2) is the softmax cross-entropy loss to optimize the learnable prompt set $\mathcal{P}$ and the classifier parameter $\phi$.
+The second term learns the key set $\mathcal{K}$ by minimizing the distances between the selected keys and the corresponding query features $q(\mathbf{x})$.
+DualPrompt: The DualPrompt method [34] supplements insertable prompts based on the L2P [35] method.
 The prompts are divided into G-Prompt (General) and E-Prompt (Expert).
@@
 The prompts in both the E-Prompt and L2P are identical, primarily serving to distinguish between different tasks.
-In contrast, the G-Prompt serves as a shared prompt across different tasks, mainly representing the shared information between tasks.
+In contrast, the G-Prompt mainly represents the shared information between tasks.
 In training, DualPrompt optimizes the G-prompt and E-prompt jointly.
-Please see [33] for details.
-3.3 CODA-Prompt
-Based on DualPrompt, CODA-Prompt [28] desires an approach that is end-to-end differentiable.
-Without selecting the top-K key-value pairs in the prompt pool, CODA-Prompt builds the prompt components: $\mathbf{p}=\sum_{m}\alpha_{m}P_{m}$, where $\mathbf{P}=\{P_{1},P_{2},\ldots,P_{M}\}\subseteq\mathbb{R}^{L_{p}\times D\times M}$ is the set of prompt components, $M$ is the length of the set.
-$\alpha_{m}$ is the weight for the component $P_{m}$, which is computed using the cosine similarity of the Query-Key.
-Unlike L2P and Dual-Prompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathbf{A}$ to process the query features: $\alpha=\gamma\left(q(\mathbf{x})\odot\mathbf{A},\mathbf{K}\right)$, where $\mathbf{A}\subseteq\mathbb{R}^{D_{k}\times M}$ are learnable parameters.
-Finally, CODA-Prompt optimizes the prompt components $\mathbf{P}$, keys $\mathbf{K}$, attention $\mathbf{A}$ and the classifier.
+CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$:
+$\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i},$ (3)
+where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key.
+Unlike L2P and DualPrompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathcal{A}=\{\mathbf{A}_{1},\ldots,\mathbf{A}_{M}\}$ for $M$ prompt components to process the query features, where $\mathbf{A}_{i}\in\mathbb{R}^{D_{k}},i=1,\ldots,M$.
+The weight is computed as $\alpha_{i}=\gamma\left(q(\mathbf{x})\odot\mathbf{A}_{i},\mathbf{k}_{i}\right)$, for $i=1,\ldots,M$.
+Finally, CODA-Prompt optimizes the prompt components $\mathcal{P}$, keys $\mathcal{K}$, attention $\mathcal{A}$ and the classifier.
 Additionally, orthogonality constraints are added to reduce interference between existing and new knowledge.
-The final total loss function is:
-$\min_{\mathbf{P},\mathbf{K},\mathbf{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathbf{P}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{K}\right)+\mathcal{L}_{\text{or}}\left(\mathbf{A}\right)\right),$ (2)
-where $\mathbf{P}$, $\mathbf{K}$, $\mathbf{A}$ refer to the prompt components and corresponding keys/attention vectors during tasks, respectively.
-The orthogonality penalty loss is $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix.
-3.4 Prompt Fine-tuning
-Based on the characteristics of large models, they possess outstanding feature extraction capabilities.
-However, due to the large model size, conventional fine-tuning approaches are not particularly effective.
-Current prompt-based continual learning methods employ a pre-trained Vision Transformer (ViT) as the backbone, with the backbone component frozen, which effectively preserves the internal hidden features of tasks.
-These methods adopt a prompt fine-tuning approach, shifting the training focus to the prompt pool.
-This approach only requires tuning a small portion of the model (the prompt embeddings) during the tasks, thereby effectively mitigating the issue of catastrophic forgetting.
-Based on the manner in which prompts are incorporated, prompt fine-tuning can be categorized into two types: prompt tuning and prefix tuning.
-Figure 2: Our teacher-student model for continual distillation learning.
-The image is processed by the query function to generate a query, which is compared with $\mathbf{K}$ in the prompt pool.
-Based on different prompt extractor methods, the output prompts are divided into $\mathbf{P_{K}}$ and $\mathbf{P_{V}}$, which are inserted into the MSA layers across multiple blocks.
-The dashed boxes represent different types of optional distillation methods.
-Feature distillation uses the internal tokens output by intermediate blocks, and the class-token is used in the classification layer for logit distillation.
-Prompt Distillation adds an additional KD-Token and KD-Classifier to establish knowledge distillation with the teacher’s classification layer.
-In this case, the student model’s prediction is influenced by both the KD-Classifier and the Student-Classifier.
-Prompt Tuning prepends the prompt to the input tokens.
-Equivalent to concatenating the prompt parameters to the input query ($\mathbf{h}_{Q}$), key ($\mathbf{h}_{K}$), and values ($\mathbf{h}_{V}$) of the MSA layer, this approach will increase the output length, where the output dimension is $\mathbb{R}^{(L+L_{p})\times D}$:
-$f_{\text{Pro-T}}(P,\mathbf{h})=\text{MSA}([P;\mathbf{h}_{Q}],[P;\mathbf{h}_{K}],[P;\mathbf{h}_{V}]).$ (3)
-Prefix Tuning divides the prompt $P$ into two parts, $P_{K}$, $P_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$, respectively:
-$f_{\text{Pre-T}}(P,\mathbf{h})=\text{MSA}(\mathbf{h}_{Q},[P_{K};\mathbf{h}_{K}],[P_{V};\mathbf{h}_{V}]).$ (4)
+The final loss function is:
+$\min_{\mathcal{P},\mathcal{K},\mathcal{A},\phi}\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)\\ +\lambda\left(\mathcal{L}_{\text{or}}\left(\mathcal{P}\right)+\mathcal{L}_{\text{or}}\left(\mathcal{K}\right)+\mathcal{L}_{\text{or}}\left(\mathcal{A}\right)\right),$ (4)
+where $\mathcal{P}$, $\mathcal{K}$, $\mathcal{A}$ refer to the prompt set and corresponding keys and attention vectors during tasks, respectively.
+The orthogonality penalty loss is defined as $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix and $I$ is the identity matrix.
+3.3 Prompting with Prefix-Tuning
+After computing the prompt $\mathbf{P}\in\mathbb{R}^{L_{p}\times D}$ for an input sample, we need to insert this prompt into the ViT backbone for prompting.
+We use the prefix-tuning method.
+It divides the prompt $\mathbf{P}$ into two parts, $\mathbf{P}_{K}$, $\mathbf{P}_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$ of the multi-head self-attention (MSA) layers, respectively:
+$f_{\text{Pre-T}}(\mathbf{P},\mathbf{h})=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}_{K};\mathbf{h}_{K}],[\mathbf{P}_{V};\mathbf{h}_{V}]).$ (5)
 This method keeps the output sequence length the same as the input.
-In this paper, since we need to insert prompts across multiple blocks, we employ the prefix method to avoid adding too many additional tokens.
-Then the prompts retrieved from the prompt pool are divided into two parts, $P_{K}$ and $P_{V}$.
-4 Continual Knowledge Distillation
-To study the CDL problem, we have designed a teacher-student model based on prompt-based continual learning methods, as illustrated in Fig.
+In the CDL setup, since we need to insert prompts across multiple blocks, we employ the prefix-tuning method to avoid adding too many additional tokens.
+4 Continual Distillation Learning
+4.1 Teacher-Student Model
+To study the CDL problem, we propose a teacher-student model setup based on prompt-based continual learning methods, as illustrated in Fig.
 2.
-Both the teacher model and the student model are some prompt-based continual learner such as L2P, DualPrompt or CODA-Prompt.
-Typically, the teacher’s backbone utilizes a larger and more accurate pretrained ViT model.
+Both the teacher model and the student model are prompt-based continual learners, such as L2P [35], DualPrompt [34], or CODA-Prompt [30].
+Typically, the teacher’s backbone utilizes a larger and more accurate pre-trained ViT model.
 The framework supports using different continual learners.
-We will study three different types of knowledge distillation within the framework: logit distillation, feature distillation and prompt distillation.
-4.1 Prompt Extractor
-To select a prompt, cosine similarity is used: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{m}\right)$ between the output of a query function $q(\mathbf{x})$ and a prompt key $\mathbf{k}_{m}$.
-Each key has a corresponding value in the prompt pool, forming key-value pairs.
-We can consider using different prompt extracting methods for different prompt-based methods.
-Then the final output prompts are divided into two parts, $P_{K}$ and $P_{V}$.
-Here, we use Prefix Tuning to insert $P_{K}$ and $P_{V}$ to the ViT backbone.
-4.2 Logit Distillation
+As illustrated in Fig.
+2, an input image $\mathbf{x}$ is first processed by the query function to generate a query, which is compared with keys $\mathcal{K}$ in the prompt pool.
+Different prompt-based continual learners select the prompts $\mathbf{P}^{\text{cl}}$ for continual learning through the Prompt Extractor, following Eq.
+(1) or Eq.
+(3).
+The prompts are divided into $\mathbf{P}^{\text{cl}}_{K}$ and $\mathbf{P}^{\text{cl}}_{V}$, which are inserted into the MSA layers across multiple blocks with prefix-tuning as in Eq.
+(5).
+This process is applied both the teacher and the student.
+Within this framework, we study and build different types of knowledge distillation in continual learning, including logit distillation and feature distillation.
+Logit distillation is one of the most classic forms of knowledge distillation.
+It aims to have a smaller student model learn the logits output of a larger or more accurate teacher model.
+In our CDL setup, we built two CDL methods based on the processing of logits: normal knowledge distillation (KD) [13] and Decoupled Knowledge Distillation (DKD) [40].
+Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model.
+In Fig.
+2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features.
+This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model.
+We built two CDL methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
+Fig. 2 illustrates the logit distillation and the feature distillation in our continual distillation learning scenario.
+Training of the teacher-student model.
+First, the images of the current task are trained on the teacher model.
+The teacher model is then used to help train the student model.
+During the student model training process, the teacher model remains frozen and is only used in the inference phase to provide soft labels (logit distillation) or internal tokens (feature distillation) to the student.
+4.2 Limitations of Existing KD Methods in CDL
+Although we can apply previous knowledge distillation methods in the CDL setting, there are inherent limitations when applying these methods to prompt-based CL.
+The primary issue arises from the query-key mechanism used in the prompt selection process within prompt-based CL methods.
+In the prompt pool, different prompt components can be interpreted as distinct feature representations.
+Through the query-key mechanism, images from different tasks can select different types of prompt feature combinations, where each image has its own preferred prompt components.
+The final prompts inserted into the backbone are either these selective prompt components (L2P, DualPrompt) or their weighted combinations (CODA-Prompt).
+In prompt-based CL, a new task may select prompt components different from prior tasks, which is an effective strategy to prevent forgetting in continual learning.
+Because each task can use its own prompts from the prompt pool.
+However, for knowledge distillation, this mechanism is not effective.
+For example, in task $A$, a subset of prompt components $\mathcal{P}_{A}$ of the student model is used for learning and knowledge distillation.
+We hypothesize that these prompt components of the student $\mathcal{P}_{A}$ inherit certain information for the teacher.
+When task $B$ comes, another subset of prompt components $\mathcal{P}_{B}$ of the student model is selected for learning and knowledge distillation.
+In the extreme case, if $\mathcal{P}_{A}\bigcap\mathcal{P}_{B}=\emptyset$, the prompts in $\mathcal{P}_{B}$ has no information about the teacher model.
+Consequently, the student needs to learn from scratch about the teacher model for task $B$.
+In knowledge distillation, the goal of the student model is to learn the overall structural knowledge of the teacher model.
+Distillation should be a global process rather than being confined independently to each task.
+4.3 Knowledge Distillation based on Prompts
+To overcome the limitations of previous KD methods in CDL, we introduce a novel knowledge distillation method named Knowledge Distillation based on Prompts (KDP), which is specifically designed for the CDL problem.
+We introduce a new type of prompts named KD prompts that are globally shared between tasks and are specifically designed for knowledge distillation as illustrated in Fig.
+2.
+Since these components are shared across all tasks, they do not suffer from the limitation mentioned before.
+Moreover, KD prompts do not rely on the query-key mechanism and do not require selection among prompt components.
+When a new task comes, knowledge is continuously transferred from the teacher to the student using the KD prompts.
+Furthermore, introducing additional KD prompts increases the number of learnable prompt components for the student model, preventing all distillation-related updates from being concentrated within the original prompt pool.
+Consequently, the prompt components of the student model are divided into two categories: the original prompts for continual learning (CL prompts) and the distillation prompts for learning from the teacher model (KD prompts).
+Similar to the CL prompts, KD prompts are inserted into the ViT backbone using the prefix-tuning method as in Eq.
+(5).
+Specifically, a KD prompt $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$ is divided into two parts: $\mathbf{P}^{\text{kd}}_{K}$, $\mathbf{P}^{\text{kd}}_{V}\in\mathbb{R}^{L_{p}/2\times D}$, which are inserted into the key $\mathbf{h}_{K}$ and value $\mathbf{h}_{V}$ at the end of each MSA layer.
+The prefix-tuning for our KDP method is defined as
+$f_{\text{Pre-T}}(\mathbf{P}^{\text{cl}},\mathbf{h},\mathbf{P}^{\text{kd}})$
+$=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}^{\text{cl}}_{K};\mathbf{h}_{K},\mathbf{P}^{\text{kd}}_{K}],[\mathbf{P}^{\text{cl}}_{V};\mathbf{h}_{V},\mathbf{P}^{\text{kd}}_{V}]).$ (6)
+For processing distilled knowledge, we adopt the distillation through attention method in DeiT [31].
+In the student model, a distillation token (the KD token in Fig.
+2) is inserted at the end of the first layer of the ViT backbone.
+After being propagated through the ViT backbone, the final output of the distillation embeddings of the KD token are connected to a separate classifier named the KD classifier.
+During training, the KD classifier is used to compute the loss function with the teacher classifier using the normal logit distillation loss function.
+That is, the logits output of the teacher model serves as the “soft labels” to the student.
+The target is the softened probability distribution of the teacher model, controlled by the temperature factor $\tau$.
+Therefore, the loss function for the distillation process is:
+$\mathcal{L}_{\text{KD}}=\tau^{2}\sum_{i}p_{i}^{\mathcal{T}}\log\left(\frac{p_{i}^{\mathcal{T}}}{p_{i}^{\mathcal{S}}}\right),$ (7)
+where $\mathcal{L}_{\text{KD}}$ is the KL divergence between between the teacher’s probability distribution $p_{i}^{\mathcal{T}}$ and the student’s probability distribution $p_{i}^{\mathcal{S}}$.
+$\tau$ is the temperature parameter.
+The final training loss function of the student model is:
+$\mathcal{L}_{\text{S}}$ $=\underbrace{(1-\alpha)\mathcal{L}(g_{\phi}^{\mathcal{S}}(f_{b}^{\mathcal{S}}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),y)}_{\text{student classification loss}}$
+$+\underbrace{\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(f^{\mathcal{S}}_{b}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),g_{\phi}^{\mathcal{T}}(f^{\mathcal{T}}_{b}(\mathbf{x})))}_{\text{student knowledge distillation loss}}$
+$+\underbrace{\lambda\mathcal{L}_{\text{pool}}}_{\text{student prompt pool loss}},$ (8)
+where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively.
+$g_{\phi}^{\mathcal{T}}$ is the teacher classifier.
+$f_{b}^{\mathcal{S}}$ and $f_{b}^{\mathcal{T}}$ represent the ViT backbones of the student and teacher models, respectively, which already include the original CL prompts.
+$\mathbf{P}^{\text{kd}}_{n}$ denotes the KD prompt inserted into the $n$-th block of the ViT backbone.
+In total, KD prompts are inserted from the first to the $n$-th block.
+$\mathcal{L}_{\text{pool}}$ is the loss function involved in extracting CL prompts from the prompt pool.
+Different prompt-based CL methods have different $\mathcal{L}_{\text{pool}}$, as shown in the latter parts of Eq.
+(2) and Eq.
+(4).
+$\alpha$ and $\lambda$ are balancing weights.
+During testing, the KD classifier and the student classifier are both used for prediction.
+Methods Split ImageNet-R Split CIFAR-100
+Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
+$\varnothing$ ViT-Small $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
+ViT-Base ViT-Small KD [13] $69.91\pm 0.62$ $7.64\pm 0.71$ $83.03\pm 0.39$ $7.24\pm 0.30$
+ViT-Base ViT-Small DKD [40] $68.92\pm 0.07$ $8.39\pm 0.36$ $82.27\pm 0.20$ $7.81\pm 0.14$
+ViT-Base ViT-Small FitNets [28] $69.87\pm 0.04$ $7.38\pm 0.36$ $81.83\pm 0.05$ $8.83\pm 0.48$
+ViT-Base ViT-Small ReviewKD [4] $70.19\pm 0.16$ $7.68\pm 0.01$ $82.20\pm 0.41$ $7.54\pm 0.03$
+ViT-Base ViT-Small DeiT [31] $70.74\pm 0.20$ $6.66\pm 0.28$ $83.79\pm 0.15$ $6.58\pm 0.06$
+ViT-Base ViT-Small KDP (ours) $\mathbf{71.92\pm 0.50}$ $\mathbf{5.61\pm 0.34}$ $\mathbf{84.31\pm 0.01}$ $\mathbf{5.63\pm 0.02}$
+$\varnothing$ ViT-Base $76.42\pm 0.17$ $4.31\pm 0.18$ $86.16\pm 0.17$ $5.63\pm 0.25$
+ViT-Large ViT-Base KD [13] $76.99\pm 0.02$ $3.81\pm 0.06$ $86.27\pm 0.05$ $5.45\pm 0.06$
+ViT-Large ViT-Base DKD [40] $76.70\pm 0.17$ $4.84\pm 0.12$ $85.42\pm 0.31$ $6.55\pm 0.16$
+ViT-Large ViT-Base FitNets [28] $74.55\pm 0.14$ $6.81\pm 0.15$ $85.95\pm 0.25$ $6.56\pm 0.02$
+ViT-Large ViT-Base ReviewKD [4] $75.72\pm 0.27$ $4.14\pm 0.05$ $86.21\pm 0.61$ $5.64\pm 0.40$
+ViT-Large ViT-Base DeiT [31] $77.83\pm 0.55$ $4.51\pm 0.03$ $86.78\pm 0.15$ $5.43\pm 0.25$
+ViT-Large ViT-Base KDP (ours) $\mathbf{78.62\pm 0.57}$ $\mathbf{3.46\pm 0.53}$ $\mathbf{87.13\pm 0.09}$ $\mathbf{5.30\pm 0.06}$
+Table 1: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on CODA-Prompt [30].
+5 Experiments
+5.1 Datasets
+We utilize the CIFAR-100 [18] and ImageNet-R [12] datasets in a class-incremental continual learning setting for our experiments.
+Following previous works, we divide the ImageNet-R and CIFAR-100 datasets into 10 tasks, where each task contains 10 classes.
+5.2 Implementation Details
+We experiment with three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34] and CODAPrompt [30].
+In L2P, the prompt pool consists of a total of 30 prompt components, and a CL prompt with a length of 20 is inserted only in the first layer of the ViT backbone.
+In DualPrompt, the prompt pool contains 10 prompt components, each CL prompt has a length of 20, with G-prompts and E-prompts placed in the first two blocks and the third to fifth blocks, respectively.
+In CODA-Prompt, the prompt pool consists of 100 prompt components.
+Each CL prompt has a length of 8, and CL prompts are inserted from the first to the fifth layer of the ViT.
+In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer.
+Our KDP method by default adopts the DeiT method [31] (including both KD token and the KD classifier) to obtain predictions.
+This is an optional component that can be added or removed.
+We evaluate the results of removing the DeiT structure from KDP in the ablation study Sec.
+5.5.
+The parameter $\alpha$ used in Eq.
+(8) to balance distillation and continual learning is set to 0.5.
+$\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
+In the experiments, all models are trained only on the data of the current task, and the distillation is based on rehearsal-free models.
+Our experiments mainly tested two teacher-student knowledge distillation pairs.
+One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6].
+For all experiments, we utilized the Adam optimizer [16].
+The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs.
+The learning rate was set to $l=0.001$.
+In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$.
+Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
+5.3 Evaluation Metrics
+Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23].
+Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Eq.
+(9), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$.
+$T$ is the total number of tasks.
+The forgetting rate (Eq.
+(10)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks.
+A higher value signifies a more negative impact of the continual learning model.
+$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i},$ (9)
+$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}.$ (10)
+5.4 Continual Distillation Results
+Methods Split ImageNet-R Split CIFAR-100
+Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
+$\varnothing$ ViT-Small + L2P [35] $63.82\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
+$\varnothing$ ViT-Small + DualPrompt [34] $65.51\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
+$\varnothing$ ViT-Small + CODA-Prompt [30] $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
+ViT-Base ViT-Small KDP (ours) + L2P [35] $68.18\pm 0.03$ $\mathbf{2.08\pm 0.28}$ $81.79\pm 0.66$ $4.31\pm 0.27$
+ViT-Base ViT-Small KDP (ours) + DualPrompt [34] $68.77\pm 0.16$ $3.13\pm 0.25$ $81.78\pm 0.17$ $\mathbf{3.63\pm 0.03}$
+ViT-Base ViT-Small KDP (ours) + CODA-Prompt [30] $\mathbf{71.92\pm 0.50}$ $5.61\pm 0.34$ $\mathbf{83.72\pm 0.08}$ $6.40\pm 0.06$
+$\varnothing$ ViT-Base + L2P [35] $73.94\pm 0.22$ $4.41\pm 0.18$ $83.02\pm 0.47$ $6.06\pm 0.47$
+$\varnothing$ ViT-Base + DualPrompt [34] $73.18\pm 0.33$ $3.45\pm 0.32$ $84.66\pm 0.87$ $5.91\pm 0.34$
+$\varnothing$ ViT-Base + CODA-Prompt [30] $76.42\pm 0.17$ $4.31\pm 0.18$ $86.16\pm 0.17$ $5.63\pm 0.25$
+ViT-Large ViT-Base KDP (ours) + L2P [35] $76.91\pm 0.40$ $\mathbf{3.15\pm 0.39}$ $86.56\pm 0.22$ $4.97\pm 0.07$
+ViT-Large ViT-Base KDP (ours) + DualPrompt [34] $76.06\pm 0.12$ $3.77\pm 0.38$ $86.92\pm 0.24$ $\mathbf{4.77\pm 0.58}$
+ViT-Large ViT-Base KDP (ours) + CODA-Prompt [30] $\mathbf{78.62\pm 0.57}$ $3.46\pm 0.53$ $\mathbf{87.13\pm 0.09}$ $5.30\pm 0.06$
+Table 2: Results of prompt distillation using KD token on the ImageNet-R dataset and the CIFAR-100 dataset with different teacher-student models and different continual learning models.
+Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner.
+The complete results for the L2P and DualPrompt methods are provided in the supplementary material.
+When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation.
+In the table, DeiT [31] refers to only using the KD token and the KD classifier without incorporating our KD prompts.
+From the table, we can see that the logit distillation methods (KD [13], DKD [40]) and feature distillation methods (FitNets [28], ReviewKD [4]) fail to address the limitations discussed in Sec.
+4.2.
+The distilled knowledge is fused into the CL prompt pool, making it difficult to transfer effectively across tasks.
+As observed in the results, although these four distillation methods can slightly improve the accuracy of the original student model, the improvements are not significant and the forgetting rate remains high.
+In many cases, the forgetting effect is even worse than that of the original student model.
+A high forgetting rate indicates that the overall accuracy improvement mainly comes from distillation benefits in the current task, while the influence of distillation on previous tasks is minimal.
+This further supports our analysis of the limitations in Sec.
+4.2.
+The DeiT method, using the distillation token and the KD classifier, separates the distilled knowledge from the CL prompt pool.
+This dual processing approach for CL and KD partially mitigates the interference between task learning and knowledge distillation using prompts.
+Our KDP method further enhances the separation and cross-task transfer of distilled knowledge through KD-Prompts, ultimately achieving the best overall performance on both datasets as shown in Table 1.
+Table 2 presents the distillation results of our KDP method on the L2P, DualPrompt and CODA-Prompt models.
+This demonstrates the generalization ability of our KDP method in different Prompt-based CL approaches.
+All three prompt-based CL methods can significantly improve their performance with our KDP method.
+The combination of CODA-Prompt and our KDP method achieves state-of-the-art (SOTA) performance, particularly benefiting CL with ViT-Small as the backbone.
+It approaches the performance of larger models while significantly reducing inference computational costs.
+5.5 Ablation Studies
+The ablation study is divided into two main aspects.
+The first aspect explores the number of block layers in which prompts are inserted and the length of each prompt.
+The second aspect examines the impact of using the KD classifier and the corresponding KD token.
+Multiple Layers & KD Prompt Length: For this study, we conducted experiments under the CODA-Prompt and L2P methods.
+The teacher model backbone is ViT-Base and the student model backbone is ViT-Small.
+We used the ImageNet-R dataset, where the number of tasks is set to 10.
+Note that a KD prompt in our model is denoted as $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$.
+In the left-hand side of Fig.
+3, we show the relationship between the length of the KD prompt, denoted as $L_{p}$, and the accuracy of the KDP model.
+In this case, the number of inserted layers of the ViT backbone is fixed at 12.
+In the right-hand side of Fig.
+3, we present the relationship between the number of inserted block layers, denoted as $n$, and the accuracy.
+In this case, $L_{p}=6$.
+From the results of the two plots, it can be observed that the length of KD prompts has no strong correlation with the final performance.
+Therefore, we select $L_{p}=6$, which performs better in CODA-Prompt.
+As the number of inserted layers $n$ increases, accuracy shows an upward trend in both methods, with a more pronounced effect in L2P.
+Based on this observation, we set the number of layers to $12$.
+Figure 3: Left: The relationship between the length of KD prompt inserted into the model and accuracy.
+Right: The relationship between the number of layers where KD prompts are inserted and accuracy.
+Teacher: Vit-Base; Student: ViT-Small; Dataset: ImageNet-R.
+KD Classifier: Our method by default adopts the KD classifier structure from DeiT [31] (including the corresponding KD token) to process the soft labels from the teacher.
+Here, we conducted an experiment to analyze the impact of using the KD classifier.
+If neither KD prompts nor the KD classifier is added, the method corresponds to the normal KD [13] in logit distillation.
+If KD prompts are removed while only adding the KD classifier with the KD token, the method corresponds to DeiT [31].
+If both components are incorporated, the method follows the default KDP setup as shown in Fig.
+2.
+Additionally, we design an experiment where only KD prompts are added, in which case the student model has only a single classifier.
+As shown in Table 3, the CODA-Prompt method is more heavily relying on the KD classifier for processing.
+The best performance is achieved when the KD prompts and the KD classifier are used together.
+However, for the L2P method, the KD classifier does not contribute significantly to improving the distillation.
+In fact, removing the KD classifier structure leads to an increase in accuracy.
+However, both CODA-Prompt and L2P depend on KD prompts to achieve better performance, which further validates the effectiveness of the KDP method.
+Model KD-Prompts KD-Classifier Avg. Acc ($\uparrow$)
+ViT-Small + CODA-Prompt [30] × × $69.91$
+ViT-Small + CODA-Prompt [30] × $70.74$
+ViT-Small + CODA-Prompt [30] × $70.21$
+ViT-Small + CODA-Prompt [30] $\mathbf{71.92}$
+ViT-Small + L2P [35] × × $63.97$
+ViT-Small + L2P [35] × $64.99$
+ViT-Small + L2P [35] × $\mathbf{69.00}$
+ViT-Small + L2P [35] $68.18$
+Table 3: Ablation study on Split ImageNet-R dataset. Comparison of different models with KD prompts and KD classifier.
+6 Conclusion and Discussion
+We introduce the problem of Continual Distillation Learning (CDL), which aims to improve prompt-based continual learning models using knowledge distillation.
+We first empirically studied logit distillation and feature distillation in the CDL setup, where three different prompt-based continual learning methods (L2P, DualPrompt and CODA-Prompt) are used.
+We found that these previous KD methods are not effective for the CDL problem.
+Therefore, we proposed a novel method, named Knowledge Distillation based on Prompts (KDP), to tackle the CDL problem.
+In KDP, a new type of prompt is introduced that is mainly designed for the distillation of knowledge.
+These additional learnable prompts significantly improve the learning performance of the student model in CDL.
+Experiments on two commonly used benchmarks for continual learning demonstrate the effectiveness of the proposed KDP method.
+Limitations. CDL models require training a teacher model and a student model jointly. The total training time and memory consumption are increased.
+Acknowledgments
+This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, and the National Science Foundation (NSF) under Grant No.
+2346528.
+Supplementary Material
+7 Why limiting the scope to prompt-based continual learning?
+Model Accuracy Prompt-based?
+iCaRL [26]-ResNet18 55.25 No
+iCaRL [26]-ResNet34 56.65 No
+LWF [19]-ResNet18 44.70 No
+LWF [19]-ResNet34 41.46 No
+L2P [35]-Tiny 60.68 Yes
+L2P [35]-Small 77.71 Yes
+L2P [35]-Base 83.02 Yes
+L2P [35]-Large 86.36 Yes
+CODA-Prompt [30]-Tiny 65.05 Yes
+CODA-Prompt [30]-Small 82.18 Yes
+CODA-Prompt [30]-Base 86.16 Yes
+CODA-Prompt [30]-Large 88.97 Yes
+Table 4: Comparison of accuracy on the CIFAR100 dataset between CL methods using different sizes of CNN backbones and CL methods using different sizes of ViT backbones.
+The number of tasks in continual learning is 10.
+This is due to the following two reasons:
+  • Prompt-based continual learning methods achieve state-of-the-art performance.
+In Table 4, we compare CODA-Prompt [30] and L2P [35] against iCaRL [26] and LWF [19] which are not prompt-based, we can clearly see that CODA-Prompt and L2P achieves better accuracy.
+  • Traditional CNN-based continual learning models do not improve or have little improvement with larger backbones.
+In Table 4, we used ResNet18 and ResNet34 for iCaRL and LWF.
+The change of backbone results in 1.4% and -3.2% for iCaRL and LWF, respectively.
+8 Logit Distillation Details
 Logit distillation is one of the most classic forms of knowledge distillation.
@@
 $\hat{p}_{i}=\frac{\exp(z_{i}/\tau)}{\sum\limits_{j\neq t}\exp(z_{j}/\tau)}.$ (17)
-4.3 Feature Distillation
+Methods Split ImageNet-R Split CIFAR-100
+Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
+$\varnothing$ ViT-Small $63.82\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
+ViT-Base ViT-Small KD [13] $63.97\pm 0.62$ $6.51\pm 0.06$ $79.64\pm 0.04$ $6.35\pm 0.02$
+ViT-Base ViT-Small DKD [40] $62.91\pm 0.27$ $6.55\pm 0.17$ $78.21\pm 0.12$ $9.13\pm 0.07$
+ViT-Base ViT-Small FitNets [28] $64.29\pm 0.09$ $6.37\pm 0.17$ $79.56\pm 0.39$ $5.89\pm 0.36$
+ViT-Base ViT-Small ReviewKD [4] $63.64\pm 0.34$ $6.36\pm 0.58$ $78.50\pm 0.39$ $8.04\pm 0.75$
+ViT-Base ViT-Small DeiT [31] $64.99\pm 0.49$ $3.83\pm 0.85$ $79.56\pm 0.07$ $6.71\pm 0.16$
+ViT-Base ViT-Small KDP (ours) $\mathbf{68.18\pm 0.03}$ $\mathbf{2.08\pm 0.28}$ $\mathbf{81.79\pm 0.66}$ $\mathbf{4.31\pm 0.27}$
+$\varnothing$ ViT-Base $73.94\pm 0.22$ $4.41\pm 0.18$ $83.02\pm 0.47$ $6.06\pm 0.47$
+ViT-Large ViT-Base KD [13] $74.12\pm 0.42$ $4.60\pm 0.55$ $85.00\pm 0.34$ $\mathbf{4.48\pm 0.51}$
+ViT-Large ViT-Base DKD [40] $74.58\pm 0.01$ $4.69\pm 0.06$ $83.29\pm 0.24$ $4.99\pm 0.17$
+ViT-Large ViT-Base FitNets [28] $70.39\pm 0.23$ $5.84\pm 0.06$ $83.60\pm 0.02$ $5.21\pm 0.71$
+ViT-Large ViT-Base ReviewKD [4] $72.17\pm 0.26$ $6.11\pm 0.08$ $83.12\pm 0.65$ $7.97\pm 0.27$
+ViT-Large ViT-Base DeiT [31] $73.99\pm 0.01$ $5.09\pm 0.02$ $84.21\pm 0.71$ $6.06\pm 0.93$
+ViT-Large ViT-Base KDP (ours) $\mathbf{76.91\pm 0.40}$ $\mathbf{3.15\pm 0.39}$ $\mathbf{86.56\pm 0.22}$ $4.97\pm 0.07$
+Table 5: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on L2P [35].
+Methods Split ImageNet-R Split CIFAR-100
+Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
+$\varnothing$ ViT-Small $65.51\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
+ViT-Base ViT-Small KD [13] $65.68\pm 0.06$ $7.26\pm 0.29$ $80.16\pm 0.54$ $5.76\pm 0.18$
+ViT-Base ViT-Small DKD [40] $65.44\pm 0.05$ $7.27\pm 0.23$ $80.44\pm 0.46$ $6.96\pm 0.44$
+ViT-Base ViT-Small FitNets [28] $66.20\pm 0.14$ $5.93\pm 0.17$ $80.70\pm 0.17$ $5.73\pm 0.21$
+ViT-Base ViT-Small ReviewKD [4] $65.69\pm 0.91$ $6.56\pm 0.46$ $80.33\pm 0.23$ $5.86\pm 0.53$
+ViT-Base ViT-Small DeiT [31] $65.82\pm 0.48$ $4.00\pm 0.19$ $80.64\pm 0.32$ $5.67\pm 0.59$
+ViT-Base ViT-Small KDP (ours) $\mathbf{68.77\pm 0.16}$ $\mathbf{3.13\pm 0.25}$ $\mathbf{81.78\pm 0.17}$ $\mathbf{3.63\pm 0.03}$
+$\varnothing$ ViT-Base $73.18\pm 0.33$ $3.45\pm 0.32$ $84.66\pm 0.87$ $5.91\pm 0.34$
+ViT-Large ViT-Base KD [13] $73.90\pm 0.14$ $\mathbf{3.31\pm 0.04}$ $84.67\pm 0.53$ $\mathbf{4.52\pm 0.55}$
+ViT-Large ViT-Base DKD [40] $75.24\pm 0.33$ $4.15\pm 0.23$ $84.93\pm 0.16$ $4.95\pm 0.12$
+ViT-Large ViT-Base FitNets [28] $71.23\pm 0.04$ $5.71\pm 0.55$ $83.12\pm 0.86$ $8.33\pm 1.36$
+ViT-Large ViT-Base ReviewKD [4] $72.19\pm 0.01$ $5.72\pm 0.20$ $84.11\pm 0.92$ $5.19\pm 0.33$
+ViT-Large ViT-Base DeiT [31] $76.03\pm 0.03$ $3.90\pm 0.01$ $85.73\pm 0.27$ $5.05\pm 0.43$
+ViT-Large ViT-Base KDP (ours) $\mathbf{76.06\pm 0.12}$ $3.77\pm 0.38$ $\mathbf{86.92\pm 0.24}$ $4.77\pm 0.58$
+Table 6: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on DualPrompt [34].
+9 Feature Distillation Details
 Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model.
@@
 It is worth noting that during knowledge distillation, the backbone of the student model remains frozen while processing the internal tokens, which is a characteristic of prompt-based continual learning methods.
-The frozen part is crucial for mitigating catastrophic forgetting during the continual learning process.
-In this paper, we study two methods based on the handling of internal tokens: FitNets [26] and Review Knowledge Distillation (ReviewKD) [4].
+In this paper, we build two methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
 FitNets [28]: This method is one of the most classic feature distillation methods, aimed at adding a distillation loss to the intermediate layers.
@@
 Therefore, in the student model, the fused features obtained at each block are passed to the next higher block to form new feature fusion.
-4.4 Prompt Distillation
-In the traditional knowledge distillation, the backbone of the student model is trainable, allowing it to learn and approach the teacher model.
-However, in the prompt-based methods, the backbone is frozen, which limits the update of intermediate features extracted by the student model, making it challenging to effectively learn knowledge from the teacher model’s features.
-Similarly, the logits are also impacted during training updates.
-In continual learning, as tasks increase, the model seeks to retain information from previous tasks.
-Prompt-based continual learning leverages a pretrained and frozen ViT model to preserve knowledge across tasks.
-However, in knowledge distillation, the student model aims to be fully trained and adjusted to approach the teacher model.
-These two objectives are conflicted with each other, which creates mutual interference between the continual learning process and the knowledge distillation process.
-Subsequently, it affects the performance of the student model.
-Here, we introduce a new concept Continual Distillation Interference (CDI) to describe this phenomenon.
-CDI is commonly observed in various continual learning methods that rely on large pretrained models as we will show in our experiments.
-To address the continual interference problem of distillation, we explored a prompt distillation method as motivated by DeiT [29].
-As shown in Fig.
-2, we introduce a learnable prompt called KD-Token in the student model, which operates alongside the class token to assist in classification.
-The KD-Token in the student model can be considered as learning knowledge from the class token in the teacher model.
-After passing through the ViT backbone, the KD token is input into a newly added KD classifier.
-The KD classifier has two roles.
-First, it is used for knowledge distillation with the teacher model.
-Second, the output of the KD classifier and the output of the student classifier are averaged to generate the final classification result.
-Therefore, the KD-Token is specifically designed to assist in knowledge distillation, while the original prompt tokens $P_{K}$ and $P_{V}$, obtained from the prompt pool, are used to facilitate the continual learning process, each serving its respective role.
-The purpose of this method is to mitigate the CDI problem in prompt-based continual learning.
-The training loss function of the student model is the same as the one in Equation (6).
-However, the KD loss is between the KD classifier and teacher classifier.
-Here, we highlight the difference by:
-$\mathcal{L}_{\text{S}}$ $=(1-\alpha)\mathcal{L}(g_{\phi}^{\mathcal{S}}(f_{b}(\mathbf{x})),y)$
-$+\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(\mathbf{F}^{\mathcal{S}}_{n}[kd]),g_{\phi}^{\mathcal{T}}(\mathbf{F}^{\mathcal{T}}_{n}[cls]))+\lambda\mathcal{L}_{\text{pool}},$ (16)
-where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively.
-$g_{\phi}^{\mathcal{T}}$ is the teacher classifier.
-$n$ is the total number of blocks in the ViT backbone.
-$\mathbf{F}^{\mathcal{S}}_{n}[kd]$ represent the KD-Token after the $n_{\text{th}}$ block of the student model.
-$\mathbf{F}^{\mathcal{T}}_{n}[cls]$ denotes the class token in the teacher model.
-The knowledge distillation loss $\mathcal{L}_{\text{KD}}$ (Equation (7)) uses KL divergence between the KD classifier and teacher classifier.
-$\mathcal{L}_{\text{pool}}$ represents the loss function on the prompt pool across different continual learning methods, and $\alpha$ and $\lambda$ are balancing weights.
-Methods Split ImageNet-R Split CIFAR-100
-Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
-$\varnothing$ ViT-Small $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
-ViT-Base ViT-Small KD [11] $69.23\pm 0.05$ $8.57\pm 0.39$ $82.39\pm 0.25$ $7.33\pm 0.21$
-ViT-Base ViT-Small DKD [39] $68.92\pm 0.07$ $8.39\pm 0.36$ $82.27\pm 0.20$ $7.81\pm 0.14$
-ViT-Base ViT-Small FitNets [26] $69.87\pm 0.04$ $7.38\pm 0.36$ $81.83\pm 0.05$ $8.83\pm 0.48$
-ViT-Base ViT-Small ReviewKD [4] $70.19\pm 0.16$ $7.68\pm 0.01$ $82.20\pm 0.41$ $7.54\pm 0.03$
-ViT-Base ViT-Small KD-Token [29] $\mathbf{70.94\pm 0.33}$ $\mathbf{6.17\pm 0.07}$ $\mathbf{83.72\pm 0.08}$ $\mathbf{6.40\pm 0.06}$
-$\varnothing$ ViT-Base $75.78\pm 0.64$ $5.70\pm 0.66$ $86.16\pm 0.17$ $5.63\pm 0.25$
-ViT-Large ViT-Base KD [11] $\mathbf{77.51\pm 0.01}$ $\mathbf{3.44\pm 0.14}$ $87.41\pm 0.05$ $5.45\pm 0.06$
-ViT-Large ViT-Base DKD [39] $76.70\pm 0.17$ $4.84\pm 0.12$ $85.42\pm 0.31$ $6.55\pm 0.16$
-ViT-Large ViT-Base FitNets [26] $74.55\pm 0.14$ $6.81\pm 0.15$ $85.95\pm 0.25$ $6.56\pm 0.02$
-ViT-Large ViT-Base ReviewKD [4] $75.72\pm 0.27$ $4.14\pm 0.05$ $86.21\pm 0.61$ $5.64\pm 0.40$
-ViT-Large ViT-Base KD-Token [29] $77.28\pm 0.29$ $4.63\pm 0.16$ $\mathbf{88.10\pm 0.33}$ $\mathbf{5.39\pm 0.19}$
-Table 1: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on CODA-Prompt [28].
-5 Experimental Results
-5.1 Experimental Benchmarks and Settings
-We utilize the CIFAR-100 [16] and ImageNet-R [10] datasets in a class-incremental continual learning setting for our experiments.
-Following previous works, we divide the ImageNet-R and CIFAR-100 into 10 tasks, where each task contains 10 classes.
-We experiment with three prompt-based continual learning models, i.e., L2P [34], DualPrompt [33] and CODAPrompt [28].
-In DualPrompt, the G-prompts are set in the first two layers with a prompt length of 5, and the E-prompts are set in layers 3 to 5.
-CODA-Prompt sets the prompt length to 8, with a total of 100 prompt components inserted into the first 5 layers.
-This setup is the same for all ViT models.
-Our experiments tested two teacher-student knowledge distillation pairs.
-One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6].
-For different knowledge distillation methods, the hyperparameter $\alpha$ for the distillation loss is set to 0.5.
-We use two metrics to evaluate the models: accuracy and forgetting rate [21].
-The accuracy refers to the average accuracy of all tasks after completing all 10 tasks.
-The forgetting rate reflects the influence of learning a new task on previously completed tasks.
-A larger forgetting signifies a more negative impact of the continual learning model.
-Please see the supplementary material for the details.
-5.2 Continual Distillation Results
-Methods Split ImageNet-R Split CIFAR-100
-Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
-$\varnothing$ ViT-Small + L2P [34] $61.95\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
-$\varnothing$ ViT-Small + DualPrompt [33] $64.27\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
-$\varnothing$ ViT-Small + CODA-Prompt [28] $67.44\pm 0.46$ $8.52\pm 0.05$ $82.18\pm 0.20$ $6.48\pm 0.48$
-ViT-Base ViT-Small KD-Token [29] + L2P [34] $65.09\pm 1.05$ $5.37\pm 0.13$ $79.78\pm 0.14$ $7.07\pm 0.52$
-ViT-Base ViT-Small KD-Token [29] + DualPrompt [33] $66.81\pm 0.42$ $\mathbf{4.55\pm 0.31}$ $81.32\pm 0.36$ $\mathbf{5.67\pm 0.59}$
-ViT-Base ViT-Small KD-Token [29] + CODA-Prompt [28] $\mathbf{70.94\pm 0.33}$ $6.17\pm 0.07$ $\mathbf{83.72\pm 0.08}$ $6.40\pm 0.06$
-$\varnothing$ ViT-Base + L2P [34] $71.59\pm 1.20$ $5.65\pm 0.69$ $83.02\pm 0.47$ $6.06\pm 0.47$
-$\varnothing$ ViT-Base + DualPrompt [33] $72.44\pm 1.00$ $3.80\pm 1.23$ $84.66\pm 0.87$ $5.91\pm 0.34$
-$\varnothing$ ViT-Base + CODA-Prompt [28] $75.78\pm 0.64$ $5.70\pm 0.66$ $86.16\pm 0.17$ $5.63\pm 0.25$
-ViT-Large ViT-Base KD-Token [29] + L2P [34] $71.89\pm 0.03$ $\mathbf{3.42\pm 0.36}$ $85.2\pm 0.28$ $\mathbf{4.58\pm 0.55}$
-ViT-Large ViT-Base KD-Token [29] + DualPrompt [33] $73.63\pm 0.26$ $3.5\pm 0.25$ $85.32\pm 0.02$ $5.64\pm 0.51$
-ViT-Large ViT-Base KD-Token [29] + CODA-Prompt [28] $\mathbf{77.28\pm 0.29}$ $4.63\pm 0.16$ $\mathbf{88.10\pm 0.33}$ $5.39\pm 0.19$
-Table 2: Results of prompt distillation using KD token on the ImageNet-R dataset and the CIFAR-100 dataset with different teacher-student models and different continual learning models.
-Table 1 summarizes the results of all the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR100 datasets using CODA-Prompt as the continual learner.
-The complete results for the L2P and DualPrompt methods are provided in the supplementary material.
-When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation.
-The student model is trained with the continual learning task only.
-We hope that the knowledge distillation methods can improve over this baseline by increasing accuracy and reducing forgetting rate.
-From the table, we observe that, prompt distillation using KD-Token achieves the best overall performance on the two datasets.
-The other knowledge distillation methods exhibit certain limitations to varying extents.
-For example, KD with ViT-Small leads to increased forgetting, and feature distillation methods such as FitNets and ReviewKD even reduce accuracy in some cases.
-Although KD, DKD, FitNets, and ReviewKD may improve continual learning performance in some situations, their effects are not significant.
-They cannot simultaneously ensure both accuracy and reduced forgetting.
-This exemplifies the continual distillation interference problem mentioned in Section  4.4.
-In contrast, the prompt distillation method using KD-Token demonstrates a more significant improvement for the student model.
-Table 2 presents the distillation results of the prompt distillation (KD-Token) method combined with L2P, DualPrompt and CODA-Prompt.
-We observe that, in all cases, the method with KD-Token distillation outperforms the original model, as reflected in the accuracy and forgetting rate metrics, validating the effectiveness of this approach.
-In addition, among the three continual learning models, CODA-Prompt achieves the best performance with or without knowledge distillation.
-Balance $\mathbf{\alpha}$ 00 $0.1$ $0.3$ $0.5$ $0.7$ $0.9$ $1.0$
-Avg. Acc ($\uparrow$) $67.44$ $68.96$ $70.06$ $\mathbf{70.94}$ $69.34$ $68.24$ $68.06$
-Forgetting ($\downarrow$) $8.52$ $7.11$ $6.46$ $\mathbf{6.17}$ $6.22$ $6.40$ $6.85$
-Table 3: Ablation study of the parameter $\alpha$ based on CODA-Prompt [28] with the Vit-Base to ViT-Small teacher-student model for prompt distillation using a KD-Token.
-5.3 Continual Distillation Interference
-The prompt distillation method shifts the distillation training process to the KD-Token and KD classifier, thereby enhancing the efficiency of the distillation.
-Based on our analysis, the freezing of the backbone results in suboptimal training outcomes for logit distillation and feature distillation.
-However, making the backbone trainable can result in a severe forgetting problem.
-To investigate the impact of adding trainable components to the backbone, we conducted a comparative experiment involving the following three models: 1) Normal KD with an additional trainable MLP layer added after the last block of the backbone (Add-MLP); 2) Prompt distillation wiht a KD-Token (Add-KD-Token); and 3) the original model without distillation (No KD).
-We distill from ViT-Base to ViT-Small using CODA-Prompt.
-After training on all tasks, we recorded the test accuracy for each task.
-The results are shown in Fig.
-3.
-It can be observed that KD with an MLP improves the distillation effectiveness for the current task and the most recent task, but results in severe forgetting for earlier tasks.
-Essentially, the objective of continual learning is to retain information from previous tasks as much as possible (model rigidity), while knowledge distillation aims for the student model to learn from the teacher model as effectively as possible (model adaptability).
-This introduces a mutual interference between continual learning and distillation, which we refer to as Continual Distillation Interference (CDI).
-The performance of Add_MLP demonstrates the existence of this problem.
-The KD-Token method attempts to address this problem by adding a KD-Token, shifting the distillation task to this prompt.
-We observe that the KD-Token method not only improves the accuracy of the current task through knowledge distillation but also maintains improved performance on previous tasks.
-5.4 Ablation Studies
-[…]
-Figure 3: After completing all tasks, the test accuracy results for each task are recorded. The CDL method uses ViT-Base to ViT-Small, with CODA-Prompt as the baseline.
-6 Conclusion and Discussion
-We introduce the problem of Continual Distillation Learning (CDL), which considers knowledge distillation in the continual learning setup.
-We emperically studied three types of knowledge distillation methods in CDL, i.e., logit distillation, feature distillation, and prompt distillation, where three different prompt-based continual learning methods (L2P, DualPrompt, and CODA-Prompt) are utilized for this study.
-We identified the presence of the Continual Distillation Interference (CDI) phenomenon in the process of CDL.
-The experimental results demonstrate that the KD-Token method outperforms the other distillation methods, consistently helping the student model improve performance in different scenarios.
-The primary purpose of this paper is to serve as an inspiration in the hope that more research on CDL will emerge in the future.
-Limitations.
-Knowledge distillation relies on a well-performing teacher model.
-Prompt-based continual learning methods can obtain teacher models by using large ViT backbones.
-However, for traditional continual learning methods that rely on CNN backbones (e.g., EWC [15] and iCaRL [24]), it is not straightforward to obtain a well-performing teacher model.
-This is why we limit our study to prompt-based continual learning.
-In addition, since each task requires training a larger teacher model, the total training time increases.
-Acknowledgments
-This work was supported in part by the DARPA Perceptually-enabled Task Guidance (PTG) Program under contract number HR00112220005, the Sony Research Award Program, and the National Science Foundation (NSF) under Grant No.
-2346528.
-Supplementary Material
-{strip}
-Methods Split ImageNet-R Split CIFAR-100
-Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
-$\varnothing$ ViT-Small $61.95\pm 0.25$ $6.52\pm 0.31$ $77.71\pm 0.49$ $7.12\pm 0.33$
-ViT-Base ViT-Small KD [11] $63.36\pm 0.04$ $6.74\pm 0.10$ $77.92\pm 0.23$ $8.30\pm 0.29$
-ViT-Base ViT-Small DKD [39] $62.91\pm 0.27$ $6.55\pm 0.17$ $78.21\pm 0.12$ $9.13\pm 0.07$
-ViT-Base ViT-Small FitNets [26] $64.29\pm 0.09$ $6.37\pm 0.17$ $79.56\pm 0.39$ $\mathbf{5.89\pm 0.36}$
-ViT-Base ViT-Small ReviewKD [4] $63.64\pm 0.34$ $6.36\pm 0.58$ $78.50\pm 0.39$ $8.04\pm 0.75$
-ViT-Base ViT-Small KD-Token [29] $\mathbf{65.09\pm 1.05}$ $\mathbf{5.37\pm 0.13}$ $\mathbf{79.78\pm 0.14}$ $7.07\pm 0.52$
-$\varnothing$ ViT-Base $71.59\pm 1.20$ $5.65\pm 0.69$ $83.02\pm 0.47$ $6.06\pm 0.47$
-ViT-Large ViT-Base KD [11] $74.12\pm 0.42$ $4.60\pm 0.55$ $80.96\pm 0.34$ $10.17\pm 0.67$
-ViT-Large ViT-Base DKD [39] $\mathbf{74.58\pm 0.01}$ $4.69\pm 0.06$ $83.29\pm 0.24$ $4.99\pm 0.17$
-ViT-Large ViT-Base FitNets [26] $70.39\pm 0.23$ $5.84\pm 0.06$ $83.60\pm 0.02$ $5.21\pm 0.71$
-ViT-Large ViT-Base ReviewKD [4] $72.17\pm 0.26$ $6.11\pm 0.08$ $83.12\pm 0.65$ $7.97\pm 0.27$
-ViT-Large ViT-Base KD-Token [29] $71.89\pm 0.03$ $\mathbf{3.42\pm 0.36}$ $\mathbf{85.20\pm 0.28}$ $\mathbf{4.58\pm 0.55}$
-Table 4: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on L2P [34].
-7 Implementation Details
-Our experiments use two metrics to evaluate the models: accuracy and forgetting rate.
-Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Equation (17), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$.
-$T$ is the total number of tasks.
-The forgetting rate (Equation (18)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks.
-A higher value signifies a more negative impact of the continual learning model.
-$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i}$ (17)
-$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}$ (18)
-For all experiments, we utilized the Adam optimizer [14].
-The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs.
-The learning rate was set to $l=0.001$.
-In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$.
-Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
 10 Continual Distillation Results
-Tables Continual Distillation Learning: An Empirical Study of Knowledge Distillation in Prompt-based Continual Learning and 5 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [34] and DualPrompt [33] as the continual learners.
+Tables 8 and 6 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [35] and DualPrompt [34] as the continual learners.
 The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners.
@@
 Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.
-Methods Split ImageNet-R Split CIFAR-100
-Teacher Student Avg. Acc ($\uparrow$) Forgetting ($\downarrow$) Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
-$\varnothing$ ViT-Small $64.27\pm 0.11$ $5.93\pm 0.03$ $79.85\pm 0.57$ $6.12\pm 0.32$
-ViT-Base ViT-Small KD [11] $65.68\pm 0.06$ $7.26\pm 0.29$ $80.16\pm 0.54$ $5.76\pm 0.18$
-ViT-Base ViT-Small DKD [39] $65.44\pm 0.05$ $7.27\pm 0.23$ $80.44\pm 0.46$ $6.96\pm 0.44$
-ViT-Base ViT-Small FitNets [26] $66.20\pm 0.14$ $5.93\pm 0.17$ $80.70\pm 0.17$ $5.73\pm 0.21$
-ViT-Base ViT-Small ReviewKD [4] $65.69\pm 0.91$ $6.56\pm 0.46$ $80.33\pm 0.23$ $5.86\pm 0.53$
-ViT-Base ViT-Small KD-Token [29] $\mathbf{66.81\pm 0.42}$ $\mathbf{4.55\pm 0.31}$ $\mathbf{81.32\pm 0.36}$ $\mathbf{5.67\pm 0.59}$
-$\varnothing$ ViT-Base $72.44\pm 1.00$ $3.80\pm 1.23$ $84.66\pm 0.87$ $5.91\pm 0.34$
-ViT-Large ViT-Base KD [11] $73.90\pm 0.14$ $\mathbf{3.31\pm 0.04}$ $83.4\pm 0.74$ $8.23\pm 1.13$
-ViT-Large ViT-Base DKD [39] $\mathbf{75.24\pm 0.33}$ $4.15\pm 0.23$ $84.93\pm 0.16$ $\mathbf{4.95\pm 0.12}$
-ViT-Large ViT-Base FitNets [26] $71.23\pm 0.04$ $5.71\pm 0.55$ $83.12\pm 0.86$ $8.33\pm 1.36$
-ViT-Large ViT-Base ReviewKD [4] $72.19\pm 0.01$ $5.72\pm 0.20$ $84.11\pm 0.92$ $5.19\pm 0.33$
-ViT-Large ViT-Base KD-Token [29] $73.63\pm 0.26$ $3.50\pm 0.25$ $\mathbf{85.32\pm 0.02}$ $5.64\pm 0.51$
-Table 5: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on DualPrompt [33].

### v4

@@
 The CDL problem is valuable to study since the use of a larger vision transformer (ViT) leads to better performance in prompt-based continual learning.
-The distillation of knowledge from a large ViT to a small ViT can improve the inference efficiency for prompt-based CL models.
+The distillation of knowledge from a large ViT to a small ViT improves the inference efficiency for prompt-based CL models.
 We empirically found that existing KD methods such as logit distillation and feature distillation cannot effectively improve the student model in the CDL setup.
-To this end, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model.
+To address this issue, we introduce a novel method named Knowledge Distillation based on Prompts (KDP), in which globally accessible prompts specifically designed for knowledge distillation are inserted into the frozen ViT backbone of the student model.
 We demonstrate that our KDP method effectively enhances the distillation performance in comparison to existing KD methods in the CDL setup.
@@
 However, larger models introduce more computation during inference.
-To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in prompt-based continual learning.
+To address this limitation, we propose a new research problem: Knowledge Distillation (KD) from large to small models in rehearsal-free prompt-based continual learning.
 We name this problem Continual Distillation Learning (CDL), which aims to improve the performance of small models by using large models as teachers in continual learning.
-We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as the iCaRL [26] model or the LWF [19] model (see Sec.
-7 in supplementary material).
+We particularly limit the scope of this work to prompt-based CL models, since we have experimentally verified that large models may not result in better performance for the traditional CNN-based CL models such as iCaRL [26] or LWF [19] (see Sec.
+A in Appendix).
 These CNN-based models update the backbones during training.
-Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem in continual learning.
-The CDL problem is different from the previous knowledge distillation setup.
-We illustrate the differences in Fig.
+Deeper and larger CNNs tend to overfit new tasks, failing to fundamentally solve the catastrophic forgetting problem.
+The CDL problem is different from the previous knowledge distillation setup as illustrated in Fig.
 1.
@@
 In general, this is a task-incremental learning process, where both the teacher and student models can only access the data for the current task at each step.
-To explore the new CDL problem, we first conducted an empirical study by applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4] and distillation token [31], to three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34], and CODA-Prompt [30].
-For each model, we consider different teacher-student configurations.
-For example, we can use ViT-Large as the teacher model and ViT-Base as the student model.
-Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool, where the selection of prompts relies on a key-query mechanism.
+It is worth noting that some traditional continual learning methods also involve knowledge distillation, but in those cases, the teacher is typically the model trained on previous tasks under the same architecture.
+In contrast, our CDL framework is based on a fundamentally different perspective: it investigates the process where, in each task, a student model learns from a teacher model with a larger ViT backbone.
+This distillation process is expected to continue throughout the entire task-incremental learning procedure.
+When applying existing knowledge distillation methods, i.e., logit distillation [13, 40], feature distillation [28, 4], to the CDL problem, we found that the results are not satisfactory.
+Since the ViT backbone is frozen in prompt-based CL, knowledge transfer from the teacher to the student can only be based on the prompt pool.
 However, for knowledge distillation, this mechanism is not effective.
 Since each task independently reselects prompt components from the prompt pool, the knowledge embedded within them from the teacher model cannot be transferred to the next task.
-Our experimental results also show that the performance improvements achieved by these knowledge distillation methods are not significant.
-Previous KD methods are not effective in the CDL setting.
-To address this limitation, we propose a novel method named Knowledge Distillation based on Prompts (KDP).
+We refer to this phenomenon as distillation information forgetting.
+To address this problem, we propose a novel method named Knowledge Distillation based on Prompts (KDP).
 KDP inserts globally accessible prompts, specifically designed for knowledge distillation, into the frozen ViT backbone of the student model.
@@
 They serve as auxiliary information to guide the learning process of the student model.
-Experimental results demonstrate that this novel approach effectively addresses the CDL problem and enhances distillation performance compared to previous KD methods.
-Furthermore, we conducted an ablation study on the total number of KD prompts inserted into the backbone.
-The results show that the introduction of KD prompts into each ViT block significantly improves the alignment of the feature structures between layers of the teacher model and the student model, leading to better performance of the student model.
-In summary, our contributions are as follows.
-  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning.
-  • We conducted an empirical study by applying different knowledge distillation methods to three prompt-based CL approaches and identify the limitations of these distillation methods in the CDL setting.
-  • To address the unique challenges of the CDL problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches.
+To enhance generalization, we conducted experiments on three representative prompt-based CL methods: L2P, DualPrompt, and CODA-Prompt.
+The results demonstrate that our KDP method effectively addresses this issue and achieves state-of-the-art performance compared to other distillation approaches.
+In summary, our contributions are as follows:
+  • We introduce the new problem of continual distillation learning (CDL) that studies knowledge distillation of a teacher model to a student model in continual learning and conduct an empirical study by applying different knowledge distillation methods to different prompt-based CL approaches.
+  • We are the first to discover and propose the distillation information forgetting problem presented in existing distillation methods and provide an analysis of its underlying causes.
+  • To address this problem, we propose a novel distillation method, KDP, which outperforms previous KD approaches in continual distillation learning.
 2 Related Work
-2.1 Continual Learning
+Continual Learning.
 The main purpose of continual learning is to build an intelligent system to solve the problem of catastrophic forgetting [24] in the case of incremental tasks.
@@
 The above methods are primarily based on CNN backbones, and a larger backbone model may not achieve better performance.
-2.2 Prompt-based Continual Learning
+Prompt-based Continual Learning.
 Inspired by the use of prompts in natural language processing [21], prompt-based continual learning methods employ large vision transformer (ViT) models [6].
@@
 CPrompt [9] introduces two main components during training: Classifier Consistency Learning (CCL) and Prompt Consistency Learning (PCL).
-In prompt-based methods, the size of the backbone matters.
-Larger backbone models achieve better performance.
+In prompt-based methods, larger backbone models achieve better performance.
 This motivates us to study the knowledge distillation problem for prompt-based continual learning methods, where we can distill knowledge from large models to small models in order to improve small models with fast inference time.
-2.3 Knowledge Distillation
-The purpose of knowledge distillation is to leverage larger models to assist smaller models to fully exploit their potential.
+Knowledge Distillation.
+The purpose of knowledge distillation is to leverage larger models to improve smaller models.
 The larger model, termed the teacher, aids in training the smaller model, termed the student.
@@
 3 Prerequisites
-3.1 Continual Learning Setting
 In continual learning, a model is required to learn a sequence of tasks, where the data from these tasks arrive on time.
@@
 In testing, the model is evaluated by testing samples from all classes.
-Figure 2: Our teacher-student model for continual distillation learning.
-The green and red dashed boxes represent different types of optional distillation methods.
-The blue dashed box represents our KDP method, where the components inside are the trainable modules to be added.
-3.2 Prompt-based Methods
 Prompt-based continual learning methods use pre-trained ViT backbones and freeze the backbones during training. Learning is shifted to trainable prompts.
@@
 The query $\mathbf{x}_{e}=q(\mathbf{x})\in\mathbb{R}^{D_{k}}$ then matches the key $\mathbf{k}_{i}$ with cosine similarity: $\gamma\left(q(\mathbf{x}),\mathbf{k}_{i}\right)$, where $\gamma(\cdot,\cdot)$ denotes the cosine similarity function.
-Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected:
-$\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M}),$ (1)
-where TopK indicates the selection function based on cosine similarity.
+Then the top $K$ prompt components from the prompt pool $\mathcal{P}$ based on the establishment of key-value pairs are selected: $\mathbf{P}_{\mathbf{x}}=\text{TopK}(\mathbf{P}_{1},\mathbf{P}_{2},\ldots,\mathbf{P}_{M})$, where TopK indicates the selection function based on cosine similarity.
 $\mathbf{P}_{\mathbf{x}}\in\mathbb{R}^{K\times L_{p}\times D}$ is the final set of prompt components extracted from the prompt pool to assist continual learning.
@@
 In training, DualPrompt optimizes the G-prompt and E-prompt jointly.
-CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$:
-$\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i},$ (3)
-where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key.
+CODA-Prompt: Instead of selecting the top $K$ key-value pairs in the prompt pool, CODA-Prompt [30] uses a weighted summation over the prompt components to compute the learnable prompt parameter for an input $\mathbf{x}$: $\mathbf{P}_{\mathbf{x}}=\sum_{i=1}^{M}\alpha_{i}\mathbf{P}_{i}$, where $\alpha_{i}$ is the weight for the prompt component $\mathbf{P}_{i}$, which is computed using the cosine similarity of the query and key.
 Unlike L2P and DualPrompt, when calculating the similarity, CODA-Prompt creates a feature-selection attention scheme $\mathcal{A}=\{\mathbf{A}_{1},\ldots,\mathbf{A}_{M}\}$ for $M$ prompt components to process the query features, where $\mathbf{A}_{i}\in\mathbb{R}^{D_{k}},i=1,\ldots,M$.
@@
 The orthogonality penalty loss is defined as $\mathcal{L}_{\text{or}}(B)=\|BB^{T}-I\|_{2}$, where $B$ represents any matrix and $I$ is the identity matrix.
-3.3 Prompting with Prefix-Tuning
-After computing the prompt $\mathbf{P}\in\mathbb{R}^{L_{p}\times D}$ for an input sample, we need to insert this prompt into the ViT backbone for prompting.
+Prompting with Prefix-Tuning: After computing the prompt $\mathbf{P}\in\mathbb{R}^{L_{p}\times D}$ for an input sample, we need to insert this prompt into the ViT backbone for prompting.
 We use the prefix-tuning method.
@@
 $f_{\text{Pre-T}}(\mathbf{P},\mathbf{h})=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}_{K};\mathbf{h}_{K}],[\mathbf{P}_{V};\mathbf{h}_{V}]).$ (3)
-This method keeps the output sequence length the same as the input.
-In the CDL setup, since we need to insert prompts across multiple blocks, we employ the prefix-tuning method to avoid adding too many additional tokens.
+This method keeps the output sequence length the same as the input. In CDL, we employ the prefix-tuning method to avoid adding too many additional tokens.
 4 Continual Distillation Learning
+Figure 2: Our teacher-student model for continual distillation learning.
+The green and red dashed boxes represent different types of optional distillation methods.
+The blue dashed box represents our KDP method, where the components inside are the trainable modules to be added.
 4.1 Teacher-Student Model
@@
 2, an input image $\mathbf{x}$ is first processed by the query function to generate a query, which is compared with keys $\mathcal{K}$ in the prompt pool.
-Different prompt-based continual learners select the prompts $\mathbf{P}^{\text{cl}}$ for continual learning through the Prompt Extractor, following Eq.
-(1) or Eq.
-(3).
+Different prompt-based continual learners select the prompts $\mathbf{P}^{\text{cl}}$ for continual learning through the Prompt Extractor, such as using TopK or weighted combination.
 The prompts are divided into $\mathbf{P}^{\text{cl}}_{K}$ and $\mathbf{P}^{\text{cl}}_{V}$, which are inserted into the MSA layers across multiple blocks with prefix-tuning as in Eq.
@@
 We built two CDL methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
-Fig. 2 illustrates the logit distillation and the feature distillation in our continual distillation learning scenario.
+Fig.
+2 illustrates the logit distillation and the feature distillation in our CDL scenario.
 Training of the teacher-student model.
@@
 The teacher model is then used to help train the student model.
-During the student model training process, the teacher model remains frozen and is only used in the inference phase to provide soft labels (logit distillation) or internal tokens (feature distillation) to the student.
-4.2 Limitations of Existing KD Methods in CDL
+During the student model training process, the teacher model remains frozen and is only used in the training phase to provide soft labels (logit distillation) or internal tokens (feature distillation) to the student.
+4.2 Distillation Information Forgetting
+Figure 3: (a) A student model may select different prompts for each new task, where teacher information is lost during distillation.
+(b) To address this limitation, one solution is to unfreeze part of the ViT, which results in a severe forgetting problem.
+(c) Our proposed solution is to introduce KD prompts, treating the ViT backbone together with the KD prompts as an overall trainable structure.
 Although we can apply previous knowledge distillation methods in the CDL setting, there are inherent limitations when applying these methods to prompt-based CL.
@@
 Distillation should be a global process rather than being confined independently to each task.
+As shown in Fig.
+3(a), for each new task, the model may re-select different CL prompts.
+Since the backbone remains frozen, only the CL prompts are updated during training.
+The distilled knowledge from the previous task is entirely contained in the CL prompts.
+Re-selecting new CL prompts leads to the loss of teacher information, a phenomenon we refer to as distillation information forgetting.
+Since existing distillation methods are not specifically tailored for CDL, the student training process, as illustrated in Fig.
+3(a), is prone to the distillation information forgetting problem.
+A potential solution is to unfreeze part of the ViT backbone (Fig.
+3(b)) and allow it to be updated during training.
+In this way, the backbone updates alongside task progression.
+However, this approach contradicts the core principle of prompt-based CL, where the backbone is frozen to preserve knowledge from previous tasks.
+Updating the backbone parameters may further exacerbate forgetting.
+We have verified this issue in Sec.
+5.2.
+Motivated by the need to address this distillation information forgetting problem, we propose a new solution by introducing KD prompts (Fig.
+3(c)).
 4.3 Knowledge Distillation based on Prompts
-To overcome the limitations of previous KD methods in CDL, we introduce a novel knowledge distillation method named Knowledge Distillation based on Prompts (KDP), which is specifically designed for the CDL problem.
+To overcome the limitations of other KD methods in CDL, we introduce a novel KD method named Knowledge Distillation based on Prompts (KDP), which is specifically designed for the CDL problem.
 We introduce a new type of prompts named KD prompts that are globally shared between tasks and are specifically designed for knowledge distillation as illustrated in Fig.
@@
 The prefix-tuning for our KDP method is defined as
-$f_{\text{Pre-T}}(\mathbf{P}^{\text{cl}},\mathbf{h},\mathbf{P}^{\text{kd}})$
-$=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}^{\text{cl}}_{K};\mathbf{h}_{K},\mathbf{P}^{\text{kd}}_{K}],[\mathbf{P}^{\text{cl}}_{V};\mathbf{h}_{V},\mathbf{P}^{\text{kd}}_{V}]).$ (6)
+$f_{\text{Pre-T}}(\mathbf{P}^{\text{cl}},\mathbf{h},\mathbf{P}^{\text{kd}})=\text{MSA}(\mathbf{h}_{Q},[\mathbf{P}^{\text{cl}}_{K};\mathbf{h}_{K},\mathbf{P}^{\text{kd}}_{K}],[\mathbf{P}^{\text{cl}}_{V};\mathbf{h}_{V},\mathbf{P}^{\text{kd}}_{V}]).$ (4)
 For processing distilled knowledge, we adopt the distillation through attention method in DeiT [31].
@@
 $\mathcal{L}_{\text{S}}$ $=\underbrace{(1-\alpha)\mathcal{L}(g_{\phi}^{\mathcal{S}}(f_{b}^{\mathcal{S}}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),y)}_{\text{student classification loss}}$
-$+\underbrace{\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(f^{\mathcal{S}}_{b}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),g_{\phi}^{\mathcal{T}}(f^{\mathcal{T}}_{b}(\mathbf{x})))}_{\text{student knowledge distillation loss}}$
-$+\underbrace{\lambda\mathcal{L}_{\text{pool}}}_{\text{student prompt pool loss}},$ (8)
+$+\underbrace{\alpha\mathcal{L}_{\text{KD}}(k_{\phi}^{\mathcal{S}}(f^{\mathcal{S}}_{b}(\mathbf{x};\mathbf{P}^{\text{kd}}_{1}:\mathbf{P}^{\text{kd}}_{n})),g_{\phi}^{\mathcal{T}}(f^{\mathcal{T}}_{b}(\mathbf{x})))}_{\text{student knowledge distillation loss}}+\underbrace{\lambda\mathcal{L}_{\text{pool}}}_{\text{student prompt pool loss}},$ (6)
 where $g_{\phi}^{\mathcal{S}}$ and $k_{\phi}^{\mathcal{S}}$ represent the student classifier and KD classifier in the student model, respectively.
@@
 $f_{b}^{\mathcal{S}}$ and $f_{b}^{\mathcal{T}}$ represent the ViT backbones of the student and teacher models, respectively, which already include the original CL prompts.
-$\mathbf{P}^{\text{kd}}_{n}$ denotes the KD prompt inserted into the $n$-th block of the ViT backbone.
+$\mathbf{P}^{\text{kd}}_{n}$ denotes the KD prompt inserted into the $n$-th block of ViT.
 In total, KD prompts are inserted from the first to the $n$-th block.
@@
 During testing, the KD classifier and the student classifier are both used for prediction.
+5 Experiments
+Datasets: We utilize the CIFAR-100 [18] and ImageNet-R [12] datasets in a class-incremental continual learning setting for our experiments.
+Following previous works, we divide the ImageNet-R and CIFAR-100 datasets into 10 tasks, where each task contains 10 classes.
+Implementation Details: We experiment with three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34] and CODAPrompt [30].
+In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer.
+Our KDP method by default adopts both KD prompts and the KD classifier to obtain predictions.
+We evaluate the results of removing the KD classifier from KDP in the ablation study Sec.
+5.2.
+The parameter $\alpha$ used in Eq.
+(6) to balance distillation and continual learning is set to 0.5.
+$\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
+In the experiments, all models are trained only on the data of the current task, and the distillation is based on rehearsal-free models.
+Our experiments mainly tested two teacher-student knowledge distillation pairs.
+One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6].
+We provide other experimental details in the Appendix B.
+Evaluation Metrics: Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23].
+Accuracy refers to the average accuracy of all tasks after completing all 10 tasks.
+The average forgetting rate, also known as backward transfer, reflects the influence of learning a new task on previously completed tasks.
+A higher value signifies a more negative impact of the continual learning model.
+We provide the details in Appendix C.
+5.1 Continual Distillation Results
 Methods Split ImageNet-R Split CIFAR-100
@@
 Table 1: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on CODA-Prompt [30].
-5 Experiments
-5.1 Datasets
-We utilize the CIFAR-100 [18] and ImageNet-R [12] datasets in a class-incremental continual learning setting for our experiments.
-Following previous works, we divide the ImageNet-R and CIFAR-100 datasets into 10 tasks, where each task contains 10 classes.
-5.2 Implementation Details
-We experiment with three prompt-based continual learning models, i.e., L2P [35], DualPrompt [34] and CODAPrompt [30].
-In L2P, the prompt pool consists of a total of 30 prompt components, and a CL prompt with a length of 20 is inserted only in the first layer of the ViT backbone.
-In DualPrompt, the prompt pool contains 10 prompt components, each CL prompt has a length of 20, with G-prompts and E-prompts placed in the first two blocks and the third to fifth blocks, respectively.
-In CODA-Prompt, the prompt pool consists of 100 prompt components.
-Each CL prompt has a length of 8, and CL prompts are inserted from the first to the fifth layer of the ViT.
-In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer.
-Our KDP method by default adopts the DeiT method [31] (including both KD token and the KD classifier) to obtain predictions.
-This is an optional component that can be added or removed.
-We evaluate the results of removing the DeiT structure from KDP in the ablation study Sec.
-5.5.
-The parameter $\alpha$ used in Eq.
-(8) to balance distillation and continual learning is set to 0.5.
-$\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
-In the experiments, all models are trained only on the data of the current task, and the distillation is based on rehearsal-free models.
-Our experiments mainly tested two teacher-student knowledge distillation pairs.
-One is distilling from ViT-Large to ViT-Base [6], and the other one is from ViT-Base to ViT-Small [6].
-For all experiments, we utilized the Adam optimizer [16].
-The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs.
-The learning rate was set to $l=0.001$.
-In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$.
-Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
-5.3 Evaluation Metrics
-Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23].
-Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Eq.
-(9), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$.
-$T$ is the total number of tasks.
-The forgetting rate (Eq.
-(10)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks.
-A higher value signifies a more negative impact of the continual learning model.
-$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i},$ (9)
-$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}.$ (10)
-5.4 Continual Distillation Results
 Methods Split ImageNet-R Split CIFAR-100
@@
 When the teacher model is $\varnothing$, it indicates that there is no knowledge distillation.
-In the table, DeiT [31] refers to only using the KD token and the KD classifier without incorporating our KD prompts.
-From the table, we can see that the logit distillation methods (KD [13], DKD [40]) and feature distillation methods (FitNets [28], ReviewKD [4]) fail to address the limitations discussed in Sec.
-4.2.
-The distilled knowledge is fused into the CL prompt pool, making it difficult to transfer effectively across tasks.
-As observed in the results, although these four distillation methods can slightly improve the accuracy of the original student model, the improvements are not significant and the forgetting rate remains high.
+As observed in the results, although these logit distillation methods (KD [13], DKD [40]) and feature distillation methods (FitNets [28] can slightly improve the accuracy of the student model, the improvements are not significant and the forgetting rate remains high.
 In many cases, the forgetting effect is even worse than that of the original student model.
@@
 4.2.
-The DeiT method, using the distillation token and the KD classifier, separates the distilled knowledge from the CL prompt pool.
+The DeiT [31] method, using the distillation token and the KD classifier, separates the distilled knowledge from the CL prompt pool.
 This dual processing approach for CL and KD partially mitigates the interference between task learning and knowledge distillation using prompts.
-Our KDP method further enhances the separation and cross-task transfer of distilled knowledge through KD-Prompts, ultimately achieving the best overall performance on both datasets as shown in Table 1.
-Table 2 presents the distillation results of our KDP method on the L2P, DualPrompt and CODA-Prompt models.
-This demonstrates the generalization ability of our KDP method in different Prompt-based CL approaches.
+Our KDP method further enhances the separation and cross-task transfer of distilled knowledge through KD-Prompts, ultimately achieving the best overall performance on both datasets.
+Table 2 presents the distillation results of our KDP method on the L2P, DualPrompt and CODA-Prompt models, which demonstrates the generalization ability of our KDP method in different Prompt-based CL approaches (the complete results can be found in Appendix Sec.
+F).
 All three prompt-based CL methods can significantly improve their performance with our KDP method.
@@
 5.2 Ablation Studies
-The ablation study is divided into two main aspects.
-The first aspect explores the number of block layers in which prompts are inserted and the length of each prompt.
-The second aspect examines the impact of using the KD classifier and the corresponding KD token.
-Multiple Layers & KD Prompt Length: For this study, we conducted experiments under the CODA-Prompt and L2P methods.
-The teacher model backbone is ViT-Base and the student model backbone is ViT-Small.
-We used the ImageNet-R dataset, where the number of tasks is set to 10.
+We perform ablation studies from multiple perspectives.
+Unless otherwise specified, all ablation experiments use CODA-Prompt as the baseline method, adopt the ViT-Base to ViT-Small setting for CDL, and are performed on the ImageNet-R dataset.
+Figure 4: Left: The relationship between the length of KD prompt and accuracy. Right: The relationship between the number of layers where KD prompts are inserted and accuracy.
+Figure 5: The effect of unfreezing the last layer in ViT. After completing all tasks, the test accuracy results for each task are recorded.
+Multiple Layers & KD Prompt Length: We conducted experiments under the CODA-Prompt and L2P methods, where the total number of tasks is set to 10.
 Note that a KD prompt in our model is denoted as $\mathbf{P}^{\text{kd}}\in\mathbb{R}^{L_{p}\times D}$.
-In the left-hand side of Fig.
-3, we show the relationship between the length of the KD prompt, denoted as $L_{p}$, and the accuracy of the KDP model.
-In this case, the number of inserted layers of the ViT backbone is fixed at 12.
-In the right-hand side of Fig.
-3, we present the relationship between the number of inserted block layers, denoted as $n$, and the accuracy.
-In this case, $L_{p}=6$.
-From the results of the two plots, it can be observed that the length of KD prompts has no strong correlation with the final performance.
-Therefore, we select $L_{p}=6$, which performs better in CODA-Prompt.
-As the number of inserted layers $n$ increases, accuracy shows an upward trend in both methods, with a more pronounced effect in L2P.
-Based on this observation, we set the number of layers to $12$.
-Figure 3: Left: The relationship between the length of KD prompt inserted into the model and accuracy.
-Right: The relationship between the number of layers where KD prompts are inserted and accuracy.
-Teacher: Vit-Base; Student: ViT-Small; Dataset: ImageNet-R.
-KD Classifier: Our method by default adopts the KD classifier structure from DeiT [31] (including the corresponding KD token) to process the soft labels from the teacher.
-Here, we conducted an experiment to analyze the impact of using the KD classifier.
+Fig.
+5 (left) shows the effect of KD prompt length $L_{p}$ on accuracy, with the number of inserted layers fixed at 12.
+Fig.
+5 (right) illustrates the impact of the number of inserted layers $n$ on accuracy, with $L_{p}=6$.
+The results show that $L_{p}$ has limited influence on performance, so we set $L_{p}=6$, which performs well in CODA-Prompt.
+In contrast, increasing $n$ improves accuracy, especially in L2P, thus we choose $n=12$ for the best performance.
+KD Classifier: Our method by default adopts the KD classifier structure (including the corresponding KD token) to process the soft labels from the teacher.
+Here, we conducted an experiment to analyze the impact of using this structure.
 If neither KD prompts nor the KD classifier is added, the method corresponds to the normal KD [13] in logit distillation.
-If KD prompts are removed while only adding the KD classifier with the KD token, the method corresponds to DeiT [31].
-If both components are incorporated, the method follows the default KDP setup as shown in Fig.
-2.
-Additionally, we design an experiment where only KD prompts are added, in which case the student model has only a single classifier.
-As shown in Table 3, the CODA-Prompt method is more heavily relying on the KD classifier for processing.
-The best performance is achieved when the KD prompts and the KD classifier are used together.
-However, for the L2P method, the KD classifier does not contribute significantly to improving the distillation.
-In fact, removing the KD classifier structure leads to an increase in accuracy.
-However, both CODA-Prompt and L2P depend on KD prompts to achieve better performance, which further validates the effectiveness of the KDP method.
-Model KD-Prompts KD-Classifier Avg. Acc ($\uparrow$)
-ViT-Small + CODA-Prompt [30] × × $69.91$
-ViT-Small + CODA-Prompt [30] × $70.74$
-ViT-Small + CODA-Prompt [30] × $70.21$
-ViT-Small + CODA-Prompt [30] $\mathbf{71.92}$
-ViT-Small + L2P [35] × × $63.97$
-ViT-Small + L2P [35] × $64.99$
-ViT-Small + L2P [35] × $\mathbf{69.00}$
-ViT-Small + L2P [35] $68.18$
-Table 3: Ablation study on Split ImageNet-R dataset. Comparison of different models with KD prompts and KD classifier.
+As shown in Table 3, the best performance is achieved when the KD prompts and the KD classifier are used together.
+Unfreezing ViT Backbone: We unfreeze the last layer (block) of the ViT backbone to allow it to participate in training.
+The experimental results in Fig.
+5 support the claim made in Sec.
+4.2 that unfreezing the part of ViT backbone does not resolve the Distillation Information Forgetting problem.
+After completing the training on the final task, we compare the per-task accuracy.
+The green curve (Unfreeze Last Layer) shows it results in overfitting to the current task and severe forgetting of earlier ones, contradicting the goal of continual learning.
+The two evaluation metrics, average accuracy dropping to 29.98% and average forgetting reaching 58.21%, indicate that unfreezing the backbone leads to poor overall performance in CDL.
+Trainable Prompt Insertion Position: Table 4 investigates the effect of prompt insert position.
+We add the same number of trainable prompts($L_{p}=6$), originally inserted according to our method, directly into the CL prompt pool.
+However, this variant is inferior compared to KDP, demonstrating that the improvement is not simply due to the increased number of trainable units.
+Instead, it shows that our KD prompts contribute to CDL by serving as global, cross-task prompts that effectively mitigate the Distillation Information Forgetting problem.
+KD-Prompts KD-Classifier Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
+$\varnothing$ $\varnothing$ $69.91$ $7.64$
+$\varnothing$ $70.74$ $6.66$
+$\varnothing$ $70.21$ $6.98$
+$\mathbf{71.92}$ $\mathbf{5.61}$
+Table 3: Comparison of the performance of the KDP method under different settings, based on whether KD prompts or the KD classifier are included.
+Position of additional prompts Avg. Acc ($\uparrow$) Forgetting ($\downarrow$)
+CL-Prompt + CODA-Prompt [30] $69.34$ $8.84$
+KD-Prompt (Ours) + CODA-Prompt [30] $\mathbf{71.92}$ $\mathbf{5.61}$
+CL-Prompt + DualPrompt [34] $65.98$ $6.82$
+KD-Prompt (Ours) + DualPrompt [34] $\mathbf{68.77}$ $\mathbf{3.13}$
+Table 4: Comparing the insertion positions of the additional prompts. CL-Prompt refers to inserting an equal number of additional trainable prompts into the CL prompt pool.
 6 Conclusion and Discussion
@@
 2346528.
-Supplementary Material
-7 Why limiting the scope to prompt-based continual learning?
+Appendix
+Appendix A Why limiting the scope to prompt-based continual learning?
 Model Accuracy Prompt-based?
@@
 The change of backbone results in 1.4% and -3.2% for iCaRL and LWF, respectively.
-8 Logit Distillation Details
+Appendix B Implementation Details
+In L2P, the prompt pool consists of a total of 30 prompt components, and a CL prompt with a length of 20 is inserted only in the first layer of the ViT backbone.
+In DualPrompt, the prompt pool contains 10 prompt components, each CL prompt has a length of 20, with G-prompts and E-prompts placed in the first two blocks and the third to fifth blocks, respectively.
+In CODA-Prompt, the prompt pool consists of 100 prompt components.
+Each CL prompt has a length of 8, and CL prompts are inserted from the first to the fifth layer of the ViT.
+In our KDP method, unless otherwise specified, all KD prompts have a length of 6 and are inserted from the first block layer to the twelfth layer.
+Our KDP method by default adopts the DeiT method [31] (including both KD token and the KD classifier) to obtain predictions.
+This is an optional component that can be added or removed.
+We evaluate the results of removing the DeiT structure from KDP in the ablation study Sec.
+5.2.
+The parameter $\alpha$ used in Eq.
+(6) to balance distillation and continual learning is set to 0.5.
+$\lambda$ for $\mathcal{L}_{\text{pool}}$ is set to 1.
+For all experiments, we utilized the Adam optimizer [16].
+The Split ImageNet-R dataset was trained for 35 epochs, while the Split CIFAR-100 dataset was trained for 20 epochs.
+The learning rate was set to $l=0.001$.
+In all loss equations, the balancing parameter was set to $\alpha=0.5$, while the temperature parameter for logit distillation was $\tau=2$.
+Training was conducted using two NVIDIA A5000 GPUs, each with 24 GB of memory.
+Appendix C Evaluation Metrics
+Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23]. Accuracy refers to the average accuracy of all tasks after completing all 10 tasks.
+Our experiments use two metrics to evaluate the models: accuracy and forgetting rate [23].
+Accuracy refers to the average accuracy of all tasks after completing all 10 tasks, defined by Eq.
+(7), where $R_{i,j}$ is the test classification accuracy of the model on task $t_{j}$ after observing the last sample from task $t_{i}$.
+$T$ is the total number of tasks.
+The forgetting rate (Eq.
+(8)), also known as backward transfer, reflects the influence of learning a new task on previously completed tasks.
+A higher value signifies a more negative impact of the continual learning model.
+$\text{ACC}=\frac{1}{T}\sum_{i=1}^{T}R_{T,i},$ (7)
+$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}R_{T,i}-R_{i,i}.$ (8)
+Appendix D Logit Distillation Details
 Logit distillation is one of the most classic forms of knowledge distillation.
@@
 $\hat{p}_{i}=\frac{\exp(z_{i}/\tau)}{\sum\limits_{j\neq t}\exp(z_{j}/\tau)}.$ (15)
+Appendix E Feature Distillation Details
+Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model.
+It can leverage richer, layer-wise information within the teacher model to guide the student model’s learning.
+The student model can benefit from an understanding of feature relationships.
+In the CDL model in Fig.
+2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features.
+This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model.
+It is worth noting that during knowledge distillation, the backbone of the student model remains frozen while processing the internal tokens, which is a characteristic of prompt-based continual learning methods.
+In this paper, we build two methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
+FitNets [28]: This method is one of the most classic feature distillation methods, aimed at adding a distillation loss to the intermediate layers.
+It uses the intermediate feature representations of the teacher model as hints to guide the student model’s learning.
+Here, we select the output of the last block of the teacher model as the hint.
+Similarly, the student model selects the output of the corresponding block for learning, constructing the feature distillation loss.
+Finally, the total loss function for the student model is
+$\mathcal{L}_{\text{S}}$ $=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{hint}}+\lambda\mathcal{L}_{\text{pool}},$ (16)
+$\mathcal{L}_{\text{hint}}$ $=\left\|f_{b-1}^{\mathcal{T}}(\mathbf{x})-F_{M}(f_{b-1}^{\mathcal{S}}(\mathbf{x}))\right\|^{2},$ (17)
+where $\mathcal{L}_{\text{hint}}$ is the feature distillation loss, calculated using the Mean Squared Error (MSE).
+$\alpha$ and $\lambda$ are used to balance the weights of loss components.
+$f_{b-1}$ indicates the feature output (internal tokens) after the last block of the ViT model.
+The student feature is transformed into the same size as the teacher feature with the mapping layer $F_{M}$, which is simply a fully-connected layer in the network.
+ReviewKD [4]: It innovates by “reviewing” multiple hidden layers from both the teacher and student models, offering a more comprehensive approach to capturing hierarchical features across the entire model.
+It reviews the multiple layers utilizing the concept of residual learning.
+For instance, the feature from $n_{th}$ block of the student is aggregated with the feature from $(n-1)_{th}$ block of the student to mimic the feature from $(n-1)_{th}$ block of the teacher.
+The total loss function of the student in ReviewKD is
+$\mathcal{L}_{\text{S}}=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{RKD}}+\lambda\mathcal{L}_{\text{pool}},$ (18)
+$\mathcal{L}_{\text{RKD}}=\mathcal{D}(\mathbf{F}^{\mathcal{S}}_{n},\mathbf{F}^{\mathcal{T}}_{n})+\sum_{j=n-1}^{1}\mathcal{D}\left(\mathcal{U}(\mathbf{F}_{j}^{\mathcal{S}},\mathbf{F}_{j+1}^{\mathcal{S}}),\mathbf{F}^{\mathcal{T}}_{j}\right),$ (19)
+where $\mathcal{L}_{\text{RKD}}$ is the reviewKD loss, and $\mathcal{D}$ is L2 distance between the student features and teacher features.
+All student features in the equations have passed through the mapping layer $F_{M}$ to make them the same dimension as the teacher features.
+$\mathbf{F}_{j}^{\mathcal{S}}$ and $\mathbf{F}_{j}^{\mathcal{T}}$ are the features output by the student model and teacher model, respectively, after passing through $j$ blocks.
+$\mathbf{F}_{j+1}^{\mathcal{S}}$ represents the fused student features at the $(j+1)_{th}$ block.
+$\mathcal{U}$ is a module used to fuse features, which performs a weighted combination of the two input features.
+$n$ is the total number of blocks in the ViT backbone.
+Therefore, in the student model, the fused features obtained at each block are passed to the next higher block to form new feature fusion.
+Appendix F Continual Distillation Results
 Methods Split ImageNet-R Split CIFAR-100
@@
 Table 7: The continual knowledge distillation results on the CIFAR-100 dataset and the ImageNet-R dataset with different teacher-student models based on DualPrompt [34].
-9 Feature Distillation Details
-Feature distillation focuses on transferring the intermediate representations from a teacher model to a student model.
-It can leverage richer, layer-wise information within the teacher model to guide the student model’s learning.
-The student model can benefit from an understanding of feature relationships.
-In the CDL model in Fig.
-2, the backbone ViT consists of multiple blocks, and we use the internal tokens outputted by each block as features.
-This allows the internal tokens of the student model to learn the information in the internal tokens of the teacher model.
-It is worth noting that during knowledge distillation, the backbone of the student model remains frozen while processing the internal tokens, which is a characteristic of prompt-based continual learning methods.
-In this paper, we build two methods based on the handling of internal tokens: FitNets [28] and Review Knowledge Distillation (ReviewKD) [4].
-FitNets [28]: This method is one of the most classic feature distillation methods, aimed at adding a distillation loss to the intermediate layers.
-It uses the intermediate feature representations of the teacher model as hints to guide the student model’s learning.
-Here, we select the output of the last block of the teacher model as the hint.
-Similarly, the student model selects the output of the corresponding block for learning, constructing the feature distillation loss.
-Finally, the total loss function for the student model is
-$\mathcal{L}_{\text{S}}$ $=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{hint}}+\lambda\mathcal{L}_{\text{pool}},$ (18)
-$\mathcal{L}_{\text{hint}}$ $=\left\|f_{b-1}^{\mathcal{T}}(\mathbf{x})-F_{M}(f_{b-1}^{\mathcal{S}}(\mathbf{x}))\right\|^{2},$ (19)
-where $\mathcal{L}_{\text{hint}}$ is the feature distillation loss, calculated using the Mean Squared Error (MSE).
-$\alpha$ and $\lambda$ are used to balance the weights of loss components.
-$f_{b-1}$ indicates the feature output (internal tokens) after the last block of the ViT model.
-The student feature is transformed into the same size as the teacher feature with the mapping layer $F_{M}$, which is simply a fully-connected layer in the network.
-ReviewKD [4]: It innovates by “reviewing” multiple hidden layers from both the teacher and student models, offering a more comprehensive approach to capturing hierarchical features across the entire model.
-It reviews the multiple layers utilizing the concept of residual learning.
-For instance, the feature from $n_{th}$ block of the student is aggregated with the feature from $(n-1)_{th}$ block of the student to mimic the feature from $(n-1)_{th}$ block of the teacher.
-The total loss function of the student in ReviewKD is
-$\mathcal{L}_{\text{S}}=\mathcal{L}(g_{\phi}(f_{b}(\mathbf{x})),y)+\alpha\mathcal{L}_{\text{RKD}}+\lambda\mathcal{L}_{\text{pool}},$ (20)
-$\mathcal{L}_{\text{RKD}}=\mathcal{D}(\mathbf{F}^{\mathcal{S}}_{n},\mathbf{F}^{\mathcal{T}}_{n})+\sum_{j=n-1}^{1}\mathcal{D}\left(\mathcal{U}(\mathbf{F}_{j}^{\mathcal{S}},\mathbf{F}_{j+1}^{\mathcal{S}}),\mathbf{F}^{\mathcal{T}}_{j}\right),$ (21)
-where $\mathcal{L}_{\text{RKD}}$ is the reviewKD loss, and $\mathcal{D}$ is L2 distance between the student features and teacher features.
-All student features in the equations have passed through the mapping layer $F_{M}$ to make them the same dimension as the teacher features.
-$\mathbf{F}_{j}^{\mathcal{S}}$ and $\mathbf{F}_{j}^{\mathcal{T}}$ are the features output by the student model and teacher model, respectively, after passing through $j$ blocks.
-$\mathbf{F}_{j+1}^{\mathcal{S}}$ represents the fused student features at the $(j+1)_{th}$ block.
-$\mathcal{U}$ is a module used to fuse features, which performs a weighted combination of the two input features.
-$n$ is the total number of blocks in the ViT backbone.
-Therefore, in the student model, the fused features obtained at each block are passed to the next higher block to form new feature fusion.
-10 Continual Distillation Results
-Tables 8 and 6 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [35] and DualPrompt [34] as the continual learners.
+Tables F and 7 present the results of the aforementioned knowledge distillation methods on the ImageNet-R and CIFAR-100 datasets, using L2P [35] and DualPrompt [34] as the continual learners.
 The results indicate that the KD-Token method consistently optimizes and improves the performance of the original student model across different datasets and various continual learners.
@@
 Additionally, on the ImageNet-R dataset, the KD-Token method attains the highest accuracy under the base-to-small scheme for both models.
+Appendix G Broader Impacts
+This work proposes KD-Prompt (KDP), a new framework for Continual Distillation Learning (CDL), which enables scalable knowledge transfer from large teacher models to compact student models across a sequence of tasks.
+In many recognition and segmentation scenarios, large vision models achieve strong performance but are impractical for deployment on resource-constrained platforms.
+Our method addresses this by continually distilling knowledge from the high-capacity teacher into a lightweight student, thus supporting continual learning while maintaining inference efficiency.
+As new tasks arrive, the teacher model acts as a continuously evolving base model, and the task-adapted student model obtained through continual distillation is used for practical deployment without catastrophic forgetting.
+This approach enables deploying compact, efficient models on latency-sensitive platforms such as mobile devices and robots, where low inference cost and fast response time are essential, while also supporting the need for continual learning.
+However, this process introduces trade-offs.
+Training large teacher models for each new task may incur substantial energy and computational costs.
+Therefore, we emphasize the need for responsible deployment, where the benefits of using smaller student models in deployment environments must be weighed against the training overhead of large-scale teacher models.
+In summary, KDP enables continual task adaptation and compression of high-performing models into compact ones suitable for real-world applications, making progress toward sustainable and practical continual learning systems.